### v2

# WorldGUI: Dynamic Testing for Comprehensive Desktop GUI Automation

Abstract
Current Graphical User Interface (GUI) agents have achieved outstanding performance in GUI element grounding. However, planning remains highly challenging, especially due to sensitivity to the initial state of the environment. Specifically, slight differences in the initial state—such as the target software not being open or the interface not being in its default state—often lead to planning errors. This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it. In this paper, we present WorldGUI, a novel GUI benchmark that designs GUI tasks with various initial states to simulate real computer-user interactions. The benchmark spans a wide range of tasks across 10 popular software applications, including PowerPoint, VSCode, and Adobe Acrobat. In addition, to address the challenges of dynamic GUI automation tasks, we propose GUI-Thinker, a holistic framework, leveraging a critique mechanism, that effectively manages the unpredictability and complexity of GUI interactions. Experimental results demonstrate that GUI-Thinker significantly outperforms Claude-3.5 (Computer Use) by 14.9% in success rate on WorldGUI tasks. This improvement underscores the effectiveness of our critical-thinking–based framework in enhancing GUI automation. The code is available at https://github.com/showlab/WorldGUI
Machine Learning, ICML
1 Introduction
Graphical User Interface (GUI) automation has become a prominent research area, driven by the need to enhance user productivity. This domain encompasses software usage, file management, office design, coding, and web browsing. Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o (OpenAI, 2023) and Claude-3.5 (Anthropic, 2024), GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user.
Figure 1: The comparison of static and dynamic testing processes. Our WorldGUI takes the first step to facilitate comprehensive GUI evaluation with various initial states. The red node represents an incorrect state.
GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition (He et al., 2016) and visual question answering (Antol et al., 2015). However, current GUI benchmarks such as OSWorld (Xie et al., 2024) and WebArena (Zhou et al., ) do not capture this dynamism. As shown on the left side of Fig. 1, most GUI benchmarks focus on initial and final states, measuring success rates but overlooking the changing initial conditions present in real GUI scenarios. These benchmarks often ignore situations where: (1) The software interface is not in its default state. (2) The agent might get user queries at any time. (3) Differences in agent robustness, where agents with the same low success rate (e.g. 20%) may vary in their ability to self-verify or self-correct, but these abilities are not measured in a static setting. As a result, these benchmarks fail to fully assess the capabilities of GUI agents.
In this paper, we take the first step toward comprehensive GUI agent testing by designing GUI tasks with various initial states. As illustrated on the right side of Fig. 1, the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default conditions, allowing tasks to start from intermediate states where users may seek assistance at any point (see Fig. 1 (b)). (2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing steps (see Fig. 1 (c)) or introducing new steps (see Fig. 1 (d)) to ensure task progression. By incorporating these elements, WorldGUI better mirrors real-world GUI environments, enabling a more accurate and thorough assessment of GUI agent capabilities.
Specifically, we present WorldGUI, a new benchmark featuring 10 desktop software applications and 315 tasks, including PowerPoint, Word, Excel, VSCode, etc. For each task, we provide a user query, an instructional video, and the corresponding project file. To ensure the task quality, we engaged four trained annotators skilled in using these applications and proficient in constructing data with scripts and agents for annotation. To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions. (See details in Sec. 3.)
In addition, we introduce a novel GUI agent framework, GUI-Thinker, which builds upon critical thinking philosophy, an aspect less emphasized in previous GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2024; Wu et al., 2024). In dynamic GUI environments, application settings may not be in default configurations. This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy. Through our analysis of real-world GUI scenarios, we identify three critical designs for a comprehensive agent: (1) Post-Planning Critique, (2) Pre-Execution Validation, and (3) Post-Action Evaluation. Specifically, GUI-Thinker comprises five core components: Planner, Planner-Critic, Step-Check, Actor, and Actor-Critic. We argue that these components are fundamental for effective GUI agents.
To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in GUI automation and propose a new benchmark WorldGUI which designs the GUI tasks with various initial states to simulate the real interactions; (2) We introduce GUI-Thinker, a comprehensive GUI framework. GUI-Thinker incorporate the thinking into the overall agent design, which provides valuable insights and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling complex tasks.
2 Related Work
Figure 2: WorldGUI: An illustration of our proposed real-world GUI benchmark. The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions. The pre-actions lead to different initial states. The key characteristic of our WorldGUI is the various initial states of the same task to stimulate the real-world testing process. The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
2.1 GUI Benchmarks
GUI benchmarks are essential for evaluating the performance and robustness of GUI agents. For web applications, WebShop (Yao et al., 2022), and WebArena (Zhou et al., ) are two text-based GUI benchmarks, the GUI information is formatted in the text style which is limited to reflect the dynamic GUI state changes. In OS environments, OSWorld (Xie et al., 2024) is a comprehensive benchmark including various operating systems with real applications. Mobile benchmarks MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) propose two GUI benchmarks of mobile applications. Windows-related benchmarks like AssistGUI (Gao et al., 2024) and WindowAgentArena (Bonatti et al., 2024) propose a list of real tasks in the Windows platform. However, these benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments. As a result, they are insufficient for comprehensively evaluating GUI agents. We summarize the differences between WorldGUI and other related benchmarks in Tab. 5.
2.2 GUI Agents
Building upon MLLMs, GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user. CogAgent (Hong et al., 2024) is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on the GUI grounding for enhancing the task performance. MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) are proposed to design the agent on mobile device. Ferret-UI (You et al., 2025) is another representative work focusing on enhancing the grounding ability in the IOS platform. AutoGLM (Liu et al., 2024) propose the GUI modes capable of learning through autonomous environmental interactions by reinforcing existing GUI models. These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction but still face limitations in handling dynamic and complicated full GUI tasks. Therefore, to enhance GUI automation in dynamic environments, we propose GUI-Thinker that improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes.
2.3 Critical Thinking in Agents
Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 (OpenAI, 2024) and Deepseek-R1 (DeepSeek-AI et al., 2025), have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks. The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency. One notable LLM-based agent framework, Reflexion (Shinn et al., 2024), demonstrates the effectiveness of self-reflection in solving complex tasks. Furthermore, Critic (Gou et al., 2023) integrates external tools into the critique process, leveraging them to improve performance. Noticing the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario. Which is not only aims to enhance the reasoning performance but also indispensable to designing the key module Actor-Critic (Konda & Tsitsiklis, 1999) to ensure task completion. A closely related work, AssistGUI (Gao et al., 2024), integrates a critical module only after the Actor module to evaluate action completion. Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor. These two modules lead to a comprehensive and fundamental GUI agent framework GUI-Thinker which will provide insights for future GUI agent design.
3 WorldGUI Benchmark
3.1 Task Formulation
GUI Automation Definition. The GUI automation task can be considered a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$ with state space $\mathcal{S}$, observation $\mathcal{O}$, action space $\mathcal{A}$, transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, and reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. In our setting, given a natural language query $q$, eg., Format the slide background with gradient fill that describes a specific task in high-level, along with an instructional video $v$ as a supplement that more detailed illustrates how to complete it, the agent first get the observation $o_{t}\in\mathcal{O}$ from the state $s_{t}\in\mathcal{S}$ in the execution environment and then generate the executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in\mathcal{S}$ and a new observation $o_{t+1}\in\mathcal{O}$. The process repeats until the task is finished or failed. The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step indicating the task completion status.
WorldGUI Task Definition. As illustrated in Fig. 2, each GUI task is paired with a user query and an instructional video. To achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case. This is accomplished through the use of pre-actions, which consist of a sequence of executable code to initialize tasks from different initial states. With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios as shown in Fig. 1.
Observation Space. The observation space $\mathcal{O}$ indicates the information of the operating system (OS) available to the agent in each state $s_{t}$. In our WorldGUI settings, we follow the previous work AssistGUI (Gao et al., 2024), encompassing two types of information: metadata $m_{t}$ from the application and screenshot $V_{t}$ of current state $s_{t}$. The metadata mainly includes the layout of panels and pop-up windows. The screenshot $V_{t}$ offers holistic visual information of the current state used for planning and action generation.
Action Space. Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others. Mouse-related actions also specify the target position in the pixel space of the observed screenshot. To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI, for controlling mouse and keyboard inputs. Each action is represented using the syntax action_type(arguments) as in Tab. 6.
3.2 Data Collection
3.2.1 Data Source
WorldGUI consists of a broad spectrum of desktop applications, which can be categorized into five main groups: (i) Office work, includes PowerPoint, Word, Excel, and Adobe Acrobat; (ii) Windows Usage, includes System Settings and File Management; (iii) Web Browsing, includes the configuration of Youtube and website operations; (iv) Coding, focus on the customization, configuration and editing of Visual Studio Code (VSCode); (v) Media Creation, uses the AI tool like Stable Diffusion to create or edit the image or videos by following user instruction.
Figure 3: GUI-Thinker. An overview of GUI-Thinker, includes five proposed components: Planner, Planner-Critic, Step-Check, Actor, and Actor-Critic. The Planner module receives the user query and an instructional video as input and generates an initial plan for the Planner-Critic process. This plan is then refined and executed step by step. Before each step is passed to the Actor module, it undergoes a Step-Check. After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes corrections if needed.
3.2.2 Pipeline of Data Construction
For each task 1) we manually collect the raw videos from the website, 2) then manually cut the raw lengthy videos into sub-clips, 3) ask annotators to manually write the user queries, 4) prepare the project files for each task for reproducibility, 5) generate the ground-truth plans by executing the scripts and agent, 6) conduct the data augmentation by varying the initial state for each task. As shown in Fig. 11, to achieve the above six steps, we invite four annotators and write the necessary scripts to help structure and format the data. Additionally, for the GT plan generation and pre-action generation, we build simple agents to obtain the data.
Raw Video Collection. We collect raw videos from the YouTube website as there are a lot of high-quality tutorials for desktop applications. For each software, we ask the annotators to watch the videos and then download them using different software.
Instruction Video Preparation. After obtaining the raw videos, we write the script codes to cut the lengthy and noisy videos into the sub-clips (30 seconds to 3 minutes) that serve as the instructional video.
User Query Generation. After obtaining the instructional videos, annotators are asked to manually write user queries corresponding to each video. For example, a user query for a task involving File Explorer might be: “Rename all the files by deleting ’class’ from their names.”
Project File Preparation. Following the AssistGUI (Gao et al., 2024), we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024). This approach guarantees that the testing process begins from a consistent state. When combined with pre-actions, it enables augmentation of the same task with various initial states.
GT Plan Generation. We write the script to accept user query $q$ and instructional video $v$ as input and generate the raw plans by agent (powered by GPT-4o). Since the raw plans are not flawless, annotators are asked to watch the videos and manually execute the tasks following the raw plans. During this process, annotators edit the plans to correct any inaccurate steps or descriptions, ultimately producing the finalized GT plans.
Pre-Actions Generation. To vary the task, we propose introducing pre-actions before the task begins. These pre-actions are created by annotators and involve corresponding scripts and agents. They are written in Python code, for example:
[…]
3.3 Evaluation
WorldGUI employs an execution-oriented evaluation approach by utilizing post-processing scripts to assess task completion. Specifically, for tasks like Office work and Web Browsing, we adopt exact matching to compare the differences between the ground-truth (GT) screenshots and the final screenshots. For tasks like File Management which would produce new folders or change the locations of files, etc. We create the shell script to check the status of files.
3.4 Data Statistics
WorldGUI compiles GUI tasks from 10 widely used applications on the Windows platform, including productivity software such as PowerPoint, Excel, and VSCode. A total of 107 meta tasks were collected from these applications, with each task being augmented 0 to 3 times based on its specific content, resulting in 208 augmented tasks. In total, WorldGUI comprises 315 tasks. See the details in Sec. C.
4 GUI-Thinker: Thinking before Doing
In this section, we introduce a new comprehensive GUI framework GUI-Thinker with a core and essential mechanism: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  (Hong et al., 2024; Cheng et al., 2024; Lin et al., 2024b; Zhang et al., 2023; Agashe et al., 2024). The GUI-Thinker includes the five fundamental but essential components as in Fig. 3 and an Interaction reasoning loop detailed in Algorithm 1. We summarize our design principles in the following:
$\bullet$ Post-Planning Critique: After the planning phase, a critique module verifies and, if necessary, self-corrects the generated plans to ensure their accuracy.
$\bullet$ Pre-Execution Validation: Before executing each subtask, a validation module determines whether the subtask should be executed. This step is crucial, as the current GUI environment may indicate that the subtask is unnecessary or requires modification to align with the current state conditions.
$\bullet$ Post-Action Evaluation: After each action execution, a mechanism evaluates whether the action was successfully completed before proceeding to the next subtask.
These critique mechanisms ensure the reliability and adaptability of GUI-Thinker in complex GUI environments.
4.1 State-Aware Planner
The State-Aware Planner processes the instructional video $v$ and user query $q$ generates an initial plan as shown in the left of Fig. 4. We use the speech recognition model Whisper (Radford et al., 2023) to translate the video $v$ into the subtitle and then send it to the MLLM for task planning. The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task. Under each $p_{i}$, there is a list of subtasks $[S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone. To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state. This additional context allows the agent to output plans that align with the actual state. For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again. By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs.
Figure 4: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic will assess the correctness of the plan and provide necessary corrections.
4.2 Planner-Critic
Post-Planning Critique. The goal of the Planner-Critic is to assess the correctness of the initial plans generated by the State-Aware Planner and provide corrections if needed. This module is designed to ensure the accuracy of the plans while leveraging the self-critique capabilities of MLLMs. As illustrated in Fig. 4, for each Initial Plan, the output consists of four components:
(1) <Flag>: Indicates whether the Initial Plan is correct.
(2) <Feedback>: Identifies the error type, categorized into one of three groups: “Wrong Steps,” “Missing Steps,” or “Redundant Steps.”
(3) <Correction>: Provide the corrected plans if the Flag indicates that the Initial Plan is incorrect.
(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons. As related works, CoT (Wei et al., 2022), GPT-o1 (OpenAI, 2024), and Deepseek-R1 (DeepSeek-AI et al., 2025) demonstrate that generating reasoning steps along with the answer would enhance the performance.
Figure 5: Step-Check. This module first checks the step completion status via an MLLM, and then based on the output navigates current task processing.
Figure 6: Actor-Critic. This module includes two parts: task verification and task correction. The design follows the verify-then-correct mechanism. We set an internal loop in this module to iteratively correct the actions.
4.3 Step-Check
Pre-Execution Validation. After the plan assessment, a navigation mechanism is crucial before sending each subtask $S_{t}=S_{j}^{i}$ at the time step $t$ to the Actor module. To address this, we designed a new module called Step-Check. Through extensive investigation, we discovered that during GUI task testing, perfect execution plans are rarely feasible due to the unpredictable nature of real application environments. Most software retains user preferences (e.g., remember the last configuration of user), meaning that when executing a specific task, the plan $p$ generated by the Planner might not align with the actual state of the software. Therefore, the model must determine whether to proceed with a subtask $S_{t}$ based on the current state (screenshot: $V_{t}$, metadata: $M_{t}$).
As illustrated in Fig. 5, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into four types:
(1) <Modify>: Indicates that the subtask should be modified or additional subtasks should be added.
(2) <Pass>: Indicates that the current subtask is unnecessary and can be skipped.
(3) <Continue>: Indicates that the subtask is valid and should be executed as planned.
(4) <Finished>: Indicates that the subtask has already been completed and requires no further action.
In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs ##\##Cannot confirm”. When this occurs, we design a Region Search module implemented by an LLM. This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region. It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the origin screenshot). The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
Method MLLM Office Win. Usage Web Coding Media Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
AssistGUI GPT-4o 26.7 26.4 34.8 21.4 33.3 20.5 36.4 27.3 42.9 21.4 26.7
Computer Use Claude-3.5 28.9 22.0 30.4 23.8 71.4 46.2 54.5 31.8 42.9 21.4 32.4
GUI-Thinker GPT-4o 44.4 31.9 34.8 21.4 47.6 38.5 45.5 27.3 57.1 50.0 36.2
GUI-Thinker Claude-3.5 57.8 41.8 43.5 23.8 76.2 48.7 63.6 40.9 57.1 57.1 47.3
Human 88.9 85.7 100.0 97.6 95.2 82.1 81.8 81.8 71.4 71.4 87.6
Table 1: Success rate (%) of different agents on WorldGUI. Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model. Meta represents the meta task, while Aug. represents the augmented task.
Method Office Win. Usage Web Coding Media Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Full Model 44.4 31.9 34.8 21.4 47.6 38.5 45.5 27.3 57.1 50.0 36.2
– w/o Planner-Critic 31.1 26.4 21.7 19.0 38.1 30.8 36.4 22.7 42.9 35.7 27.9
– w/o Step-Check 31.1 30.8 21.7 19.0 33.3 35.9 45.5 22.7 28.6 28.6 29.2
– w/o Actor-Critic 15.6 11.0 4.3 4.8 28.6 23.1 0.0 4.5 28.6 28.6 13.3
Table 2: Success rate (%) of our GUI-Thinker (Full Model) with the ablation of different critical modules.
4.4 Actor
The goal of the Actor is to translate natural language subtask $S_{t}$ into executable code $C_{t}$. Using an MLLM as the backbone model, the Actor processes metadata $m_{t}$ and screenshot $V_{t}$ as GUI context to generate precise executable actions, such as click(100, 200). Additionally, it leverages the history of previous actions as memory to aid in generating subsequent actions. The generated actions will executed in the environment, and then the new screenshot $V_{t+1}$ and metadata $m_{t+1}$ will be captured for the next processing.
4.5 Actor-Critic
[…]
MLLM PPT Word Excel VSCode Acrobat Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Claude-3.5 54.5 47.8 54.5 45.0 58.3 30.4 63.6 40.9 63.6 44.0 47.3
GPT-4o 45.5 39.1 36.4 25.0 50.0 34.8 54.5 27.2 45.5 32.0 36.1
Gemini-2.0-Flash 36.4 34.8 27.3 20.0 41.7 34.8 54.5 36.3 18.2 24.2 32.0
Table 3: Success rate (%) of our GUI-Thinker with the ablation on MLLM.
Method PPT Word Excel Overall
Full Model 41.2 29.0 40.0 37.0
w/o Inst. Video 35.3 22.6 22.2 29.0
Table 4: Success rate (%) of our GUI-Thinker with the ablation of Instructional Video (Inst. Video).
5 Experiments
Implementation Details. We implement the MLLM in our GUI-Thinker by using GPT-4o (OpenAI, 2023) (gpt-4o-2024-08-06) by default. For the computer mouse and keyboard control, we use the Python library PyAutoGUI. Following the AssistGUI (Gao et al., 2024), we use the GUI parser to obtain the position information of elements, e.g., buttons, icons, and text. We use some vision foundation models such as Google OCR to extract the text. By default, we use the center coordinates to represent the location of each element. All the testing is under the same screenshot resolution (1920 $\times$ 1080). In all experiments, we set the max trials of the Actor-Critic to 3 for light interaction costs. For the total trials of each task, we set it to $12\times N+$ 1, where $N$ is the length of subtasks $S_{i}$ in plan $p$. In most cases, we set the trials to 21.
Evaluation. Given that our WorldGUI includes 315 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results.
Metric. Following the previous works of OSworld and AssistGUI, we use Success Rate (SR) as the metric.
Baselines. We compare our GUI-Thinker with two strong approaches: AssistGUI (Gao et al., 2024) and Cluade 3.5 (Computer Use) (Anthropic, 2024). AssistGUI is a prominent agent framework designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query. We implement it by increasing the MLLM to GPT-4o for better performance. Cluade 3.5 (Computer Use) is the strongest MLLM model specially designed for autonomous computer use. We use the open-source implementation OOTB (Hu et al., 2024) as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison.
5.1 Main Results on WorldGUI
As shown in Tab. 1, our GUI-Thinker (Claude-3.5) outperforms previous agents, achieving the highest success rate. Compared to AssistGUI (GPT-4o), GUI-Thinker (GPT-4o) demonstrates superior performance in Office, Web, and Coding GUI tasks. This demonstrates that the critical modules are essential for effectively handling complex tasks. However, the overall success rate (SR) remains low, indicating that GUI tasks in WorldGUI are still challenging, as reflected in the Human SR of 87.6%. When comparing the Meta and augmented (Aug.) tasks, we observe that all agents perform better on the Meta tasks but struggle with the Aug. tasks, suggesting that current agents are not yet effective in handling tasks with varying states. Analyzing the results of Computer Use (Claude-3.5), we find that it excels in Web and Coding GUI tasks, achieving a 71.4% SR on Web tasks and 54.5% SR on Coding tasks, likely due to its design being optimized for computer usage. These results significantly surpass those of GUI-Thinker (GPT-4o), which lacks the corresponding GUI automation capabilities. Additionally, by equipping GUI-Thinker with Claude-3.5 as the base model, our agent achieves comparable performance in Web and Coding tasks. For the Office task, Computer Use (Claude-3.5) struggles, likely due to the dense domain-specific icons and buttons in these applications, which pose challenges for perception. By incorporating Claude-3.5 as the MLLM in our GUI-Thinker, GUI-Thinker (Claude-3.5) achieves twice the SR (57.8% vs. 28.9%) on the Office task (Meta), demonstrating the effectiveness of our agent framework.
5.2 Ablation Study of GUI-Thinker
Impact of Critical Modules. In Tab. 2, we investigate the effectiveness of our proposed critical modules: Planner-Critic, Step-Check, and Actor-Critic. We use GUI-Thinker (GPT-4o) as the Full Model to reduce financial costs. The results show that removing any of these modules leads to a decline in performance across all task groups, highlighting their importance. Planner-Critic plays a crucial role in plan correction, as its removal leads to a relative performance drop across most tasks, particularly in Office (44.4 $\rightarrow$ 31.1), Web (47.6 $\rightarrow$ 38.1), and Coding (45.5 $\rightarrow$ 36.4). This indicates that effective plan refinement is essential for successfully completing GUI tasks. Step-Check is particularly important for Web and Media tasks, where removing it results in a notable decline (Web: 47.6 $\rightarrow$ 33.3, Media: 57.1 $\rightarrow$ 28.6). This suggests that some web-based tasks require additional step modifications to improve execution accuracy. Actor-Critic is the most critical module, as its removal results in the sharpest performance drop across all task groups (Overall SR: 36.2 $\rightarrow$ 13.3). We observe that many subtasks are difficult to execute correctly on the first attempt and heavily rely on the Actor-Critic for additional action correction. The most drastic declines are seen in Coding (45.5 $\rightarrow$ 0.0), Windows Usage (34.8 $\rightarrow$ 4.3), and Office (44.4 $\rightarrow$ 15.6), indicating that these tasks are more rely on the Actor-Critic module. Overall, these results demonstrate that all three modules contribute significantly to the performance, with Actor-Critic being the most impactful.
Impact of Different MLLM. We also investigate the impact of using different MLLMs in our framework. As shown in Tab. 3, we compare three MLLMs across five popular applications. Claude-3.5 consistently achieves the highest performance on all five applications, particularly excelling in VSCode and Acrobat tasks, where it attains an SR of over 60%. GPT-4o demonstrates stable performance across these applications, while Gemini-2.0 underperforms compared to the other two MLLMs.
Impact of Instructional Video. In Tab. 4, we study the impact of removing the instructional video by modifying the prompt to include only the user query for generating the initial plan. On the three Office applications, we observe a significant performance decline, as these tasks rely more heavily on additional domain knowledge for successful planning. In contrast, the MLLM performs relatively well on Win. Usage tasks, such as Settings and File Management, where it has more inherent familiarity.
Conclusion
In this paper, we take the first step toward comprehensive GUI agent evaluation by introducing a new benchmark, WorldGUI. In addition to the standard static testing processes, we incorporate dynamic testing procedures to ensure that WorldGUI effectively captures the complexity and dynamism of real-world GUI environments. Furthermore, to enhance GUI automation, we propose a novel agent framework, GUI-Thinker, built upon a critical thinking philosophy and comprising five core components. This framework enables the agent to dynamically identify uncommon states and adjust its plans or actions accordingly. Finally, we evaluate the latest computer-using agent, Claude-3.5, using our WorldGUI benchmark, demonstrating the effectiveness of GUI-Thinker across a variety of GUI tasks.
Impact Statement
This paper introduces WorldGUI and GUI-Thinker, a benchmark and agent framework designed to advance the field of GUI automation. Our work has several important societal and research implications. Firstly, WorldGUI represents a significant step toward comprehensive GUI agent evaluation by incorporating diverse initial states for each GUI task. This approach better reflects real-world conditions, inspiring future GUI-related benchmarks to emphasize realistic GUI testing scenarios in their designs. Secondly, we propose a holistic GUI agent framework, GUI-Thinker, which integrates five essential components and an iterative reasoning loop. By embedding critical thinking into its module design, GUI-Thinker enhances adaptability in complex and dynamic GUI environments, allowing agents to handle unpredictable interface changes more effectively. Our work lays a strong foundation for future research in adaptive, reasoning-driven GUI automation, with potential applications in software testing, accessibility tools, and intelligent user assistance systems.
Appendix A Comparison with other benchmarks
In this section, we summarize the data statistics of other benchmarks in Tab. 5 as the comparison. We can observe that the main difference of our WorldGUI is we incorporate various initial states to stimulate the real GUI automation while other benchmarks are not considered.
Benchmark Softwares Tasks Platform Task Focus Inst. Video GT Plan Various States
WebArena (Zhou et al., ) 6 812 Web Website Usage
VisualWebArena (Koh et al., 2024) 3 910 Web Website Usage
Mobile-Eval (Wang et al., 2024) 10 30 Android OS APP Usage
APPAgent (Zhang et al., 2023) 10 50 Android OS APP Usage
OSWorld (Xie et al., 2024) 10 369 Ubuntu+Win. Software Usage
AssistGUI (Gao et al., 2024) 9 100 Windows Productivity Software Usage
VideoGUI (Lin et al., 2024a) 10 86 Win. + Web Productivity Software Usage
WindowAgentArena (Bonatti et al., 2024) 11 154 Windows SoftWare Usage
WorldGUI 10 315 Win. + Web Productivity Software Usage
Table 5: Comparison with related benmarks. WorldGUI is a unique benchmark that has the various states for each task to stimulate the real-world agent-computer interactions.
Why Instructional Video? The rationale for using instructional videos is that, in some cases, completing a task requires extensive configuration, making it impractical to provide all necessary information within the input query. As illustrated in Fig. 7, our user query is “Generate a photo of a girl with short brown hair with EasyNegative to improve the quality”, which does not specify how to improve the quality. The instructional video will serve as contextual information to guide the agent to complete that.
Figure 7: An example of using Stable Diffusion for media creation in WorldGUI. The left shows the interface of using stable diffusion to create the photo. The right shows the user query and GT plan. If no instructional video, it is hard to include all the settings
Appendix B Details of Actor Space
In this section, we detail the action space used in our WorldGUI. Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others. Mouse-related actions also specify the target position in the pixel space of the observed screenshot. To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI, for controlling mouse and keyboard inputs. Each action is represented using the syntax action_type(arguments) as in Tab. 6.
Action Type Example
Mouse Movement moveTo(120, 200)
Mouse Clicks click(200, 300)
Keyboard Type write(’classes’)
Hotkey hotkey(’ctrl’, ’a’)
Scrolling scroll(-100)
Drag dragTo(120, 220, 2)
Mouse Down and Up mouseDown(); mouseUp()
Press Keys press(’delete’)
Key Down and Up keyDown(’shift’)
Table 6: The action types and its example in WorldGUI.
Appendix C Data
(1) Annotators. In this work, we have four annotators: A, B, C, and D. The team comprises one PhD student, one Master’s student, and two undergraduate students. Prior to annotation, all annotators receive training on using the applications in WorldGUI to ensure high-quality annotations. For the 10 desktop applications, we divide the software into four parts, assigning each part to a different annotator. For the human tests presented in Tab. 1, the annotators demonstrate tasks on software that they did not annotate. As shown in Tab. 5, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.
Annotators Annotation Phase Human Test Phase
A PowerPoint, Word, Excel VSCode, Stable Diffusion, Web
B Adobe Acrobat, Stable Diffusion Excel, Settings
C Settings, Web PowerPoint, File Explorer, Youtube
D VSCode, File Explorer Word, Adobe Acrobat
Table 7: This table shows the software used during the annotation and human testing phases by different annotators.
(2) Data Construction Pipeline. For each task 1) we manually collect the raw videos from the website, 2) then manually cut the raw lengthy videos into sub-clips, 3) ask annotators to manually write the user queries, 4) prepare the project files for each task for reproducibility, 5) generate the ground-truth plans by executing the scripts and agent, 6) conduct the data augmentation by varying the initial state for each task. As shown in Fig. 11, to achieve the above six steps, we invite four annotators and write the necessary scripts to help structure and format the data. Additionally, for the GT plan generation and pre-action generation, we build simple agents to obtain the data.
(3) Project File and Enviroment Configuration. In this paper, we utilize project files to ensure the reproducibility of task executions. This approach is simpler and more cost-effective compared to using virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024). The types of project files for each application are illustrated in Tab. 8. To enable cross-platform usage, such as controlling the computer via a phone or other systems, we implement a frontend and backend system. The frontend receives user queries, captures screenshots and metadata, and then sends this information to the backend for API calls and the invocation of local vision foundation models. Cross-platform support allows us to deploy autonomous agents in a simple and cost-effective manner.
(4) Creating Augmented Tasks. In our study, to stimulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks. Specifically, we propose pre-actions before executing the task. The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state. This approach addresses scenarios where users may seek AI assistance because they are unable to complete a task. For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu. If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure. 2) Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings. This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps. We illustrate one example in Fig. 8. Here, the meta task and augmented task, have the same user query and instructional video and it will ideally have the same final state. We additionally provide more examples about augmenting the meta task in Fig. 9.
Figure 8: An example of augmenting one GUI task with manually aug the initial state and then using the execution scripts and corresponding agents to obtain the pre-action for each augmented case.
Figure 9: We present the examples of conducting the augmentations from the meta task.
(5) Data Statistic. WorldGUI comprises 315 GUI tasks sourced from 10 widely used Windows applications, including PowerPoint, Excel, Word, Adobe Acrobat, VSCode, File Explorer, Settings, Web, Youtube, and Stable Diffusion. Here we provide the details about the task activities and project file for each application in Tab. 8. We collected 107 meta tasks from these applications, each of which was augmented 0 to 3 times based on its specific content, resulting in 208 augmented tasks. Fig. 10 presents the task distribution, provides query examples, and illustrates the distribution of meta and augmented tasks across different applications.
Applications All Task Task Activities Project File Type
PowerPoint 34 Change the content style and layout; Design new effects. project.pptx
Word 31 Formatting the content style and layout project.docx
Excel 35 Table formatting; Data management and processing project.xlsx
Adobe Acrobat 36 Automatic add electric signature; Document management project.pdf
VSCode 33 Code editing; Software configuration vscode.exe
Settings 39 Advanced personalized and safety settings; ms-Settings
File Explorer 26 File management: Add, delete, rename, and move files. explorer.exe
Web 32 Web operation web browser + URL
Youtube 28 Video and account configurations web browser + UR
Stable Diffusion 21 Media creation with complicated effects Local Run Platform
Table 8: This table shows all tasks, task activities, and project file of the desktop applications used in WorldGUI.
Figure 10: Distribution of collect tasks, selected queries, and task amount of WorldGUI. We have gathered tasks across 10 desktop applications, focusing on the use of productivity software as well as fundamental computer operations and settings.
Figure 11: Pipeline of Data Construction. Human: Represents the annotators. Code: Refers to the scripts (e.g., Python Code) utilized to achieve the goal. Agent: We design an agent built upon the MLLMs to achieve the goal.
Appendix D GUI-Thinker Reasoning Loop Algorithm
In this section, we provide the details of our reasoning loop algorithm in Algorithm 1.
Algorithm 1 GUI-Thinker Reasoning Loop Algorithm
  Input: State $s$, Action Code $C$, Screenshot $V$, Metadata $m$, Current subtask $S$, Critic_count $z$
  Generate task plan $p$ with Planner and Planner-Critic
  Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of $p$.
  Initial $s_{0}=\textless Continue\textgreater$ >
  while $S_{t}$ is not end and $t<$ max trials do
     Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
     Obtain state $s_{t}$ by running Step-Check.
     if $s_{t}=\textless Next\textgreater$ > then
        Go to the next task $S_{t+1}=next(S_{t})$
     end if
     Check potential modification of subtask $S_{t}$
     Obtain action code $C_{t}$ by running Actor; Execute the action code $C_{t}$ in the Env.; Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
     Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ > (For each subtask, the first step is finished, then execute the actor-critic process)
     Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
     Running Actor-Critic and obtain the state $s_{t}$
     if $s_{t}=\textless Next\textgreater$ > then
        Go to the next task $S_{t+1}=next(S_{t})$.
     end if
     while $s_{t}=\textless Critic\textgreater$ > and $z<$ max critique trials do
        Running Actor-Critic and obtain the state $s_{t}$ and corrected action code $C_{t}$
        if $s_{t}=\textless Next\textgreater$ > then
           Go to the next task $S_{t+1}=next(S_{t})$.
        end if
        Execute the action code $C_{t}$ in the Env.; Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
        Set $C_{t}=$ None; $z=z+1$
     end while
     Go to the next task $S_{t+1}=next(S_{t})$
     $t=t+1$
  end while
Appendix E Qualitative Results
(1) In Fig. 12, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
Figure 12: We show one successful prediction of our GUI-Thinker.
Figure 13: We show two examples of using GUI Parser to obtain the element position information.
(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Fig. 14, Fig. 15, and Fig. 16. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
Figure 14: An example of using Planner-Critic to correct the plan.
Figure 15: Two examples of using Step-Check to check the subtask status.
Figure 16: An example of using Actor-Critic to correct the actions.
(3) We also highlight some common errors encountered. 1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown in the left of Fig. 17, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Fig. 17, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Fig. 18. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task success.
Figure 17: We display some common errors.
Figure 18: We display some common errors

### v3

@@
-WorldGUI: Dynamic Testing for Comprehensive Desktop GUI Automation
+WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point
 Abstract
-Current Graphical User Interface (GUI) agents have achieved outstanding performance in GUI element grounding.
-However, planning remains highly challenging, especially due to sensitivity to the initial state of the environment.
-Specifically, slight differences in the initial state—such as the target software not being open or the interface not being in its default state—often lead to planning errors.
+GUI agents have achieved outstanding performance in GUI element grounding.
+However, planning remains highly challenging, especially due to the sensitivity to the initial state of the environment.
+Specifically, slight differences in the initial state-such as the target software not being open or the interface not being in its default state, often lead to planning errors.
 This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it.
-In this paper, we present WorldGUI, a novel GUI benchmark that designs GUI tasks with various initial states to simulate real computer-user interactions.
-The benchmark spans a wide range of tasks across 10 popular software applications, including PowerPoint, VSCode, and Adobe Acrobat.
-In addition, to address the challenges of dynamic GUI automation tasks, we propose GUI-Thinker, a holistic framework, leveraging a critique mechanism, that effectively manages the unpredictability and complexity of GUI interactions.
-Experimental results demonstrate that GUI-Thinker significantly outperforms Claude-3.5 (Computer Use) by 14.9% in success rate on WorldGUI tasks.
-This improvement underscores the effectiveness of our critical-thinking–based framework in enhancing GUI automation.
-The code is available at https://github.com/showlab/WorldGUI
-Machine Learning, ICML
+To address this gap, we introduce WorldGUI, a comprehensive GUI benchmark containing tasks across ten widely used desktop and web applications (e.g., PowerPoint, VSCode, Acrobat), each instantiated with diverse initial states to simulate authentic human–computer interactions.
+Complementing this, we propose WorldGUI-Agent, a universal framework that unifies three core modules: Planner-Critic for high-level plan refinement, Step-Check for intermediate verification, and Actor-Critic for action-level optimization to proactively detect and correct errors.
+Experimental evaluation shows that WorldGUI-Agent outperforms the outstanding existing model (Claude-3.5 Computer Use) by 12.4% in success rate on WorldGUI, and achieves a 31.2% overall success rate on WindowsAgentArena, surpassing the prior state-of-the-art by 11.7%.
+Our analysis further reveals that dynamic augmentation tasks and desktop environments pose substantial hurdles, underscoring the necessity of adaptive planning and feedback-driven execution for advancing real-world GUI automation.
+The code and data are available at https://github.com/showlab/WorldGUI.
+00footnotetext: Corresponding author.
+Figure 1: Software taxonomy of WorldGUI and the performance comparison of GUI agents.
+The left shows 5 main groups and 10 software in our WorldGUI.
+The right shows that WorldGUI-Agent surpasses previous SOTA GUI agents significantly.
 1 Introduction
@@
 This domain encompasses software usage, file management, office design, coding, and web browsing.
-Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o (OpenAI, 2023) and Claude-3.5 (Anthropic, 2024), GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user.
-Figure 1: The comparison of static and dynamic testing processes.
-Our WorldGUI takes the first step to facilitate comprehensive GUI evaluation with various initial states.
-The red node represents an incorrect state.
-GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition (He et al., 2016) and visual question answering (Antol et al., 2015).
-However, current GUI benchmarks such as OSWorld (Xie et al., 2024) and WebArena (Zhou et al., ) do not capture this dynamism.
-As shown on the left side of Fig.
-1, most GUI benchmarks focus on initial and final states, measuring success rates but overlooking the changing initial conditions present in real GUI scenarios.
+Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o gpt4o and Claude-3.5 claude3.5 , GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to enhance productivity efficiency.
+GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition he2016deep and visual question answering antol2015vqa ; VQAv2 .
+However, current online GUI benchmarks such as WebArena webarena , WebVoyager webvoyager , and WindowsAgentArena windowsagentarena do not capture this dynamism.
+Currently, most GUI benchmarks OSWorld ; windowsagentarena ; assistgui ; webarena ; visualwebarena ; webvoyager focus on initial and final states, measuring success rates but overlooking the state variety in real GUI scenarios.
 These benchmarks often ignore situations where: (1) The software interface is not in its default state.
-(2) The agent might get user queries at any time.
-(3) Differences in agent robustness, where agents with the same low success rate (e.g.
-20%) may vary in their ability to self-verify or self-correct, but these abilities are not measured in a static setting.
-As a result, these benchmarks fail to fully assess the capabilities of GUI agents.
-In this paper, we take the first step toward comprehensive GUI agent testing by designing GUI tasks with various initial states.
-As illustrated on the right side of Fig.
-1, the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default conditions, allowing tasks to start from intermediate states where users may seek assistance at any point (see Fig.
-1 (b)).
-(2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing steps (see Fig.
-1 (c)) or introducing new steps (see Fig.
-1 (d)) to ensure task progression.
-By incorporating these elements, WorldGUI better mirrors real-world GUI environments, enabling a more accurate and thorough assessment of GUI agent capabilities.
-Specifically, we present WorldGUI, a new benchmark featuring 10 desktop software applications and 315 tasks, including PowerPoint, Word, Excel, VSCode, etc.
-For each task, we provide a user query, an instructional video, and the corresponding project file.
-To ensure the task quality, we engaged four trained annotators skilled in using these applications and proficient in constructing data with scripts and agents for annotation.
-To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions.
-(See details in Sec.
-3.)
-In addition, we introduce a novel GUI agent framework, GUI-Thinker, which builds upon critical thinking philosophy, an aspect less emphasized in previous GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2024; Wu et al., 2024).
-In dynamic GUI environments, application settings may not be in default configurations.
-This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy.
-Through our analysis of real-world GUI scenarios, we identify three critical designs for a comprehensive agent: (1) Post-Planning Critique, (2) Pre-Execution Validation, and (3) Post-Action Evaluation.
-Specifically, GUI-Thinker comprises five core components: Planner, Planner-Critic, Step-Check, Actor, and Actor-Critic.
-We argue that these components are fundamental for effective GUI agents.
-To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in GUI automation and propose a new benchmark WorldGUI which designs the GUI tasks with various initial states to simulate the real interactions; (2) We introduce GUI-Thinker, a comprehensive GUI framework.
-GUI-Thinker incorporate the thinking into the overall agent design, which provides valuable insights and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling complex tasks.
-2 Related Work
-Figure 2: WorldGUI: An illustration of our proposed real-world GUI benchmark.
+(2) The human-computer interactions may start from the intermediate state of a specific task.
+(3) Differences in agent robustness, where agents with the same low success rate (e.g., 20%) may vary in their ability to self-reflection, but these abilities cannot be measured in a static setting.
+As a result, these benchmarks fail to comprehensively assess the GUI agents.
+Figure 2: WorldGUI.
 The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions.
@@
 The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
-2.1 GUI Benchmarks
-GUI benchmarks are essential for evaluating the performance and robustness of GUI agents.
-For web applications, WebShop (Yao et al., 2022), and WebArena (Zhou et al., ) are two text-based GUI benchmarks, the GUI information is formatted in the text style which is limited to reflect the dynamic GUI state changes.
-In OS environments, OSWorld (Xie et al., 2024) is a comprehensive benchmark including various operating systems with real applications.
-Mobile benchmarks MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) propose two GUI benchmarks of mobile applications.
-Windows-related benchmarks like AssistGUI (Gao et al., 2024) and WindowAgentArena (Bonatti et al., 2024) propose a list of real tasks in the Windows platform.
-However, these benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments.
-As a result, they are insufficient for comprehensively evaluating GUI agents.
-We summarize the differences between WorldGUI and other related benchmarks in Tab.
-5.
-2.2 GUI Agents
-Building upon MLLMs, GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user.
-CogAgent (Hong et al., 2024) is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on the GUI grounding for enhancing the task performance.
-MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) are proposed to design the agent on mobile device.
-Ferret-UI (You et al., 2025) is another representative work focusing on enhancing the grounding ability in the IOS platform.
-AutoGLM (Liu et al., 2024) propose the GUI modes capable of learning through autonomous environmental interactions by reinforcing existing GUI models.
-These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction but still face limitations in handling dynamic and complicated full GUI tasks.
-Therefore, to enhance GUI automation in dynamic environments, we propose GUI-Thinker that improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes.
-2.3 Critical Thinking in Agents
-Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 (OpenAI, 2024) and Deepseek-R1 (DeepSeek-AI et al., 2025), have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks.
-The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency.
-One notable LLM-based agent framework, Reflexion (Shinn et al., 2024), demonstrates the effectiveness of self-reflection in solving complex tasks.
-Furthermore, Critic (Gou et al., 2023) integrates external tools into the critique process, leveraging them to improve performance.
-Noticing the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario.
-Which is not only aims to enhance the reasoning performance but also indispensable to designing the key module Actor-Critic (Konda & Tsitsiklis, 1999) to ensure task completion.
-A closely related work, AssistGUI (Gao et al., 2024), integrates a critical module only after the Actor module to evaluate action completion.
-Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor.
-These two modules lead to a comprehensive and fundamental GUI agent framework GUI-Thinker which will provide insights for future GUI agent design.
+In this paper, we take the first step toward comprehensive GUI evaluation by designing GUI tasks with various initial states.
+We consider that the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default initial conditions, allowing tasks to start from intermediate states where users may seek assistance at any point.
+(2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing plans or introducing new steps to ensure task execution.
+By incorporating these situations into the benchmark design, WorldGUI better mirrors real-world GUI interactions, enabling a more accurate and thorough assessment of GUI agent capabilities.
+Specifically, WorldGUI embraces 10 widely-used desktop applications with 611 tasks in total.
+For each task, we create a user query, an instructional video, and the corresponding project file.
+We engaged four trained annotators skilled in using these applications for annotation.
+To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions.
+In addition, we introduce a new GUI agent framework, WorldGUI-Agent, which builds upon critical thinking design principle, an aspect less emphasized in previous GUI agents hong2024cogagent ; cheng2024seeclick ; autowebglm ; AgentS ; osatlas .
+In dynamic GUI environments, application settings may not be in default configurations.
+This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy.
+Through our analysis of real-world GUI scenarios, we identify three design principles for GUI agents: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation.
+We argue that these components are fundamental and universal for GUI agents.
+To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in the online GUI testing and propose a new benchmark called WorldGUI; (2) We introduce WorldGUI-Agent, a fundamental and universal GUI framework that incorporates critical thinking into the overall agent design, providing valuable insight and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling GUI tasks (see Figure 1).
 2 WorldGUI Benchmark
+Table 1: Comparison with other interactive GUI benchmarks.
+WorldGUI is a unique benchmark that embraces diverse initial states and better reflects the authentic interactions in GUI scenarios.
+Env?: Indicates whether an environment is required to be deployed.
+Benchmarks Softwares Tasks Platform Env? Inst. Video? GT Plan Diverse Contextual
+Init. State? Variability?
+WebArena webarena 6 812 Web Yes
+VisualWebArena visualwebarena 3 910 Web Yes
+WebVoyager webvoyager 15 643 Web Yes
+AutoDroid AutoDroid 13 158 Android OS Yes
+AndroidWorld androidworld 20 116 Android OS Yes
+AgentStudio zheng2025agentstudio 9 205 Desktop + Web Yes
+Mobile-Eval wang2024mobile 10 30 Android OS Yes
+APPAgent zhang2023appagent 10 50 Android OS Yes
+OSWorld OSWorld 10 369 Desktop Yes
+AssistGUI assistgui 9 100 Windows No
+WindowAgentArena windowsagentarena 11 154 Windows Yes
+WorldGUI 10 611 Win. + Web No
 2.1 Task Formulation
@@
 In our setting, given a natural language query $q$, eg., Format the slide background with gradient fill that describes a specific task in high-level, along with an instructional video $v$ as a supplement that more detailed illustrates how to complete it, the agent first get the observation $o_{t}\in\mathcal{O}$ from the state $s_{t}\in\mathcal{S}$ in the execution environment and then generate the executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in\mathcal{S}$ and a new observation $o_{t+1}\in\mathcal{O}$.
-The process repeats until the task is finished or failed.
-The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step indicating the task completion status.
+The process repeats until the task is finished or fails.
+The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step ,indicating the task completion status.
 WorldGUI Task Definition.
-As illustrated in Fig.
-2, each GUI task is paired with a user query and an instructional video.
-To achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case.
+As illustrated in Figure  2, to achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case.
 This is accomplished through the use of pre-actions, which consist of a sequence of executable code to initialize tasks from different initial states.
-With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios as shown in Fig.
-1.
+With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios.
+We additionally summarize the differences between WorldGUI and other close interactive benchmarks in Table 1.
 Observation Space.
 The observation space $\mathcal{O}$ indicates the information of the operating system (OS) available to the agent in each state $s_{t}$.
-In our WorldGUI settings, we follow the previous work AssistGUI (Gao et al., 2024), encompassing two types of information: metadata $m_{t}$ from the application and screenshot $V_{t}$ of current state $s_{t}$.
-The metadata mainly includes the layout of panels and pop-up windows.
+In this paper, we follow the previous work of AssistGUI assistgui , encompassing two types of information: metadata $m_{t}$ from the application and screenshot $V_{t}$ of the current state $s_{t}$.
+The metadata mainly includes the layout of panels and UI trees.
 The screenshot $V_{t}$ offers holistic visual information of the current state used for planning and action generation.
-Action Space.
-Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others.
-Mouse-related actions also specify the target position in the pixel space of the observed screenshot.
-To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI, for controlling mouse and keyboard inputs.
-Each action is represented using the syntax action_type(arguments) as in Tab.
-6.
-3.2 Data Collection
-3.2.1 Data Source
-WorldGUI consists of a broad spectrum of desktop applications, which can be categorized into five main groups: (i) Office work, includes PowerPoint, Word, Excel, and Adobe Acrobat; (ii) Windows Usage, includes System Settings and File Management; (iii) Web Browsing, includes the configuration of Youtube and website operations; (iv) Coding, focus on the customization, configuration and editing of Visual Studio Code (VSCode); (v) Media Creation, uses the AI tool like Stable Diffusion to create or edit the image or videos by following user instruction.
-Figure 3: GUI-Thinker.
-An overview of GUI-Thinker, includes five proposed components: Planner, Planner-Critic, Step-Check, Actor, and Actor-Critic.
-The Planner module receives the user query and an instructional video as input and generates an initial plan for the Planner-Critic process.
-This plan is then refined and executed step by step.
-Before each step is passed to the Actor module, it undergoes a Step-Check.
-After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes corrections if needed.
-3.2.2 Pipeline of Data Construction
-For each task 1) we manually collect the raw videos from the website, 2) then manually cut the raw lengthy videos into sub-clips, 3) ask annotators to manually write the user queries, 4) prepare the project files for each task for reproducibility, 5) generate the ground-truth plans by executing the scripts and agent, 6) conduct the data augmentation by varying the initial state for each task.
-As shown in Fig.
-11, to achieve the above six steps, we invite four annotators and write the necessary scripts to help structure and format the data.
-Additionally, for the GT plan generation and pre-action generation, we build simple agents to obtain the data.
-Raw Video Collection.
-We collect raw videos from the YouTube website as there are a lot of high-quality tutorials for desktop applications.
-For each software, we ask the annotators to watch the videos and then download them using different software.
-Instruction Video Preparation.
-After obtaining the raw videos, we write the script codes to cut the lengthy and noisy videos into the sub-clips (30 seconds to 3 minutes) that serve as the instructional video.
-User Query Generation.
-After obtaining the instructional videos, annotators are asked to manually write user queries corresponding to each video.
-For example, a user query for a task involving File Explorer might be: “Rename all the files by deleting ’class’ from their names.”
-Project File Preparation.
-Following the AssistGUI (Gao et al., 2024), we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024).
-This approach guarantees that the testing process begins from a consistent state.
-When combined with pre-actions, it enables augmentation of the same task with various initial states.
-GT Plan Generation.
-We write the script to accept user query $q$ and instructional video $v$ as input and generate the raw plans by agent (powered by GPT-4o).
-Since the raw plans are not flawless, annotators are asked to watch the videos and manually execute the tasks following the raw plans.
-During this process, annotators edit the plans to correct any inaccurate steps or descriptions, ultimately producing the finalized GT plans.
-Pre-Actions Generation.
-To vary the task, we propose introducing pre-actions before the task begins.
-These pre-actions are created by annotators and involve corresponding scripts and agents.
-They are written in Python code, for example:
-[…]
-3.3 Evaluation
-WorldGUI employs an execution-oriented evaluation approach by utilizing post-processing scripts to assess task completion.
-Specifically, for tasks like Office work and Web Browsing, we adopt exact matching to compare the differences between the ground-truth (GT) screenshots and the final screenshots.
-For tasks like File Management which would produce new folders or change the locations of files, etc.
-We create the shell script to check the status of files.
-3.4 Data Statistics
-WorldGUI compiles GUI tasks from 10 widely used applications on the Windows platform, including productivity software such as PowerPoint, Excel, and VSCode.
-A total of 107 meta tasks were collected from these applications, with each task being augmented 0 to 3 times based on its specific content, resulting in 208 augmented tasks.
-In total, WorldGUI comprises 315 tasks.
-See the details in Sec.
-C.
-4 GUI-Thinker: Thinking before Doing
-In this section, we introduce a new comprehensive GUI framework GUI-Thinker with a core and essential mechanism: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  (Hong et al., 2024; Cheng et al., 2024; Lin et al., 2024b; Zhang et al., 2023; Agashe et al., 2024).
-The GUI-Thinker includes the five fundamental but essential components as in Fig.
-3 and an Interaction reasoning loop detailed in Algorithm 1.
-We summarize our design principles in the following:
-$\bullet$ Post-Planning Critique: After the planning phase, a critique module verifies and, if necessary, self-corrects the generated plans to ensure their accuracy.
-$\bullet$ Pre-Execution Validation: Before executing each subtask, a validation module determines whether the subtask should be executed.
-This step is crucial, as the current GUI environment may indicate that the subtask is unnecessary or requires modification to align with the current state conditions.
-$\bullet$ Post-Action Evaluation: After each action execution, a mechanism evaluates whether the action was successfully completed before proceeding to the next subtask.
-These critique mechanisms ensure the reliability and adaptability of GUI-Thinker in complex GUI environments.
-4.1 State-Aware Planner
-The State-Aware Planner processes the instructional video $v$ and user query $q$ generates an initial plan as shown in the left of Fig.
-4.
-We use the speech recognition model Whisper (Radford et al., 2023) to translate the video $v$ into the subtitle and then send it to the MLLM for task planning.
-The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task.
-Under each $p_{i}$, there is a list of subtasks $[S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone.
-To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state.
-This additional context allows the agent to output plans that align with the actual state.
-For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again.
-By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs.
-Figure 4: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic will assess the correctness of the plan and provide necessary corrections.
-4.2 Planner-Critic
-Post-Planning Critique.
-The goal of the Planner-Critic is to assess the correctness of the initial plans generated by the State-Aware Planner and provide corrections if needed.
-This module is designed to ensure the accuracy of the plans while leveraging the self-critique capabilities of MLLMs.
-As illustrated in Fig.
-4, for each Initial Plan, the output consists of four components:
-(1) <Flag>: Indicates whether the Initial Plan is correct.
-(2) <Feedback>: Identifies the error type, categorized into one of three groups: “Wrong Steps,” “Missing Steps,” or “Redundant Steps.”
-(3) <Correction>: Provide the corrected plans if the Flag indicates that the Initial Plan is incorrect.
-(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons.
-As related works, CoT (Wei et al., 2022), GPT-o1 (OpenAI, 2024), and Deepseek-R1 (DeepSeek-AI et al., 2025) demonstrate that generating reasoning steps along with the answer would enhance the performance.
-Figure 5: Step-Check. This module first checks the step completion status via an MLLM, and then based on the output navigates current task processing.
-Figure 6: Actor-Critic.
-This module includes two parts: task verification and task correction.
-The design follows the verify-then-correct mechanism.
-We set an internal loop in this module to iteratively correct the actions.
-4.3 Step-Check
-Pre-Execution Validation.
-After the plan assessment, a navigation mechanism is crucial before sending each subtask $S_{t}=S_{j}^{i}$ at the time step $t$ to the Actor module.
-To address this, we designed a new module called Step-Check.
-Through extensive investigation, we discovered that during GUI task testing, perfect execution plans are rarely feasible due to the unpredictable nature of real application environments.
-Most software retains user preferences (e.g., remember the last configuration of user), meaning that when executing a specific task, the plan $p$ generated by the Planner might not align with the actual state of the software.
-Therefore, the model must determine whether to proceed with a subtask $S_{t}$ based on the current state (screenshot: $V_{t}$, metadata: $M_{t}$).
-As illustrated in Fig. 5, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into four types:
-(1) <Modify>: Indicates that the subtask should be modified or additional subtasks should be added.
-(2) <Pass>: Indicates that the current subtask is unnecessary and can be skipped.
-(3) <Continue>: Indicates that the subtask is valid and should be executed as planned.
-(4) <Finished>: Indicates that the subtask has already been completed and requires no further action.
-In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs ##\##Cannot confirm”.
-When this occurs, we design a Region Search module implemented by an LLM.
-This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region.
-It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the origin screenshot).
-The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
-Method MLLM Office Win. Usage Web Coding Media Overall
-Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
-AssistGUI GPT-4o 26.7 26.4 34.8 21.4 33.3 20.5 36.4 27.3 42.9 21.4 26.7
-Computer Use Claude-3.5 28.9 22.0 30.4 23.8 71.4 46.2 54.5 31.8 42.9 21.4 32.4
-GUI-Thinker GPT-4o 44.4 31.9 34.8 21.4 47.6 38.5 45.5 27.3 57.1 50.0 36.2
-GUI-Thinker Claude-3.5 57.8 41.8 43.5 23.8 76.2 48.7 63.6 40.9 57.1 57.1 47.3
-Human 88.9 85.7 100.0 97.6 95.2 82.1 81.8 81.8 71.4 71.4 87.6
-Table 1: Success rate (%) of different agents on WorldGUI.
-Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model.
-Meta represents the meta task, while Aug.
-represents the augmented task.
-Method Office Win. Usage Web Coding Media Overall
-Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
-Full Model 44.4 31.9 34.8 21.4 47.6 38.5 45.5 27.3 57.1 50.0 36.2
-– w/o Planner-Critic 31.1 26.4 21.7 19.0 38.1 30.8 36.4 22.7 42.9 35.7 27.9
-– w/o Step-Check 31.1 30.8 21.7 19.0 33.3 35.9 45.5 22.7 28.6 28.6 29.2
-– w/o Actor-Critic 15.6 11.0 4.3 4.8 28.6 23.1 0.0 4.5 28.6 28.6 13.3
-Table 2: Success rate (%) of our GUI-Thinker (Full Model) with the ablation of different critical modules.
-4.4 Actor
-The goal of the Actor is to translate natural language subtask $S_{t}$ into executable code $C_{t}$.
-Using an MLLM as the backbone model, the Actor processes metadata $m_{t}$ and screenshot $V_{t}$ as GUI context to generate precise executable actions, such as click(100, 200).
-Additionally, it leverages the history of previous actions as memory to aid in generating subsequent actions.
-The generated actions will executed in the environment, and then the new screenshot $V_{t+1}$ and metadata $m_{t+1}$ will be captured for the next processing.
-4.5 Actor-Critic
-[…]
-MLLM PPT Word Excel VSCode Acrobat Overall
-Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
-Claude-3.5 54.5 47.8 54.5 45.0 58.3 30.4 63.6 40.9 63.6 44.0 47.3
-GPT-4o 45.5 39.1 36.4 25.0 50.0 34.8 54.5 27.2 45.5 32.0 36.1
-Gemini-2.0-Flash 36.4 34.8 27.3 20.0 41.7 34.8 54.5 36.3 18.2 24.2 32.0
-Table 3: Success rate (%) of our GUI-Thinker with the ablation on MLLM.
-Method PPT Word Excel Overall
-Full Model 41.2 29.0 40.0 37.0
-w/o Inst. Video 35.3 22.6 22.2 29.0
-Table 4: Success rate (%) of our GUI-Thinker with the ablation of Instructional Video (Inst. Video).
-5 Experiments
-Implementation Details.
-We implement the MLLM in our GUI-Thinker by using GPT-4o (OpenAI, 2023) (gpt-4o-2024-08-06) by default.
-For the computer mouse and keyboard control, we use the Python library PyAutoGUI.
-Following the AssistGUI (Gao et al., 2024), we use the GUI parser to obtain the position information of elements, e.g., buttons, icons, and text.
-We use some vision foundation models such as Google OCR to extract the text.
-By default, we use the center coordinates to represent the location of each element.
-All the testing is under the same screenshot resolution (1920 $\times$ 1080).
-In all experiments, we set the max trials of the Actor-Critic to 3 for light interaction costs.
-For the total trials of each task, we set it to $12\times N+$ 1, where $N$ is the length of subtasks $S_{i}$ in plan $p$.
-In most cases, we set the trials to 21.
-Evaluation. Given that our WorldGUI includes 315 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results.
-Metric. Following the previous works of OSworld and AssistGUI, we use Success Rate (SR) as the metric.
-Baselines.
-We compare our GUI-Thinker with two strong approaches: AssistGUI (Gao et al., 2024) and Cluade 3.5 (Computer Use) (Anthropic, 2024).
-AssistGUI is a prominent agent framework designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query.
-We implement it by increasing the MLLM to GPT-4o for better performance.
-Cluade 3.5 (Computer Use) is the strongest MLLM model specially designed for autonomous computer use.
-We use the open-source implementation OOTB (Hu et al., 2024) as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison.
-5.1 Main Results on WorldGUI
-As shown in Tab.
-1, our GUI-Thinker (Claude-3.5) outperforms previous agents, achieving the highest success rate.
-Compared to AssistGUI (GPT-4o), GUI-Thinker (GPT-4o) demonstrates superior performance in Office, Web, and Coding GUI tasks.
-This demonstrates that the critical modules are essential for effectively handling complex tasks.
-However, the overall success rate (SR) remains low, indicating that GUI tasks in WorldGUI are still challenging, as reflected in the Human SR of 87.6%.
-When comparing the Meta and augmented (Aug.) tasks, we observe that all agents perform better on the Meta tasks but struggle with the Aug.
-tasks, suggesting that current agents are not yet effective in handling tasks with varying states.
-Analyzing the results of Computer Use (Claude-3.5), we find that it excels in Web and Coding GUI tasks, achieving a 71.4% SR on Web tasks and 54.5% SR on Coding tasks, likely due to its design being optimized for computer usage.
-These results significantly surpass those of GUI-Thinker (GPT-4o), which lacks the corresponding GUI automation capabilities.
-Additionally, by equipping GUI-Thinker with Claude-3.5 as the base model, our agent achieves comparable performance in Web and Coding tasks.
-For the Office task, Computer Use (Claude-3.5) struggles, likely due to the dense domain-specific icons and buttons in these applications, which pose challenges for perception.
-By incorporating Claude-3.5 as the MLLM in our GUI-Thinker, GUI-Thinker (Claude-3.5) achieves twice the SR (57.8% vs.
-28.9%) on the Office task (Meta), demonstrating the effectiveness of our agent framework.
-5.2 Ablation Study of GUI-Thinker
-Impact of Critical Modules.
-In Tab.
-2, we investigate the effectiveness of our proposed critical modules: Planner-Critic, Step-Check, and Actor-Critic.
-We use GUI-Thinker (GPT-4o) as the Full Model to reduce financial costs.
-The results show that removing any of these modules leads to a decline in performance across all task groups, highlighting their importance.
-Planner-Critic plays a crucial role in plan correction, as its removal leads to a relative performance drop across most tasks, particularly in Office (44.4 $\rightarrow$ 31.1), Web (47.6 $\rightarrow$ 38.1), and Coding (45.5 $\rightarrow$ 36.4).
-This indicates that effective plan refinement is essential for successfully completing GUI tasks.
-Step-Check is particularly important for Web and Media tasks, where removing it results in a notable decline (Web: 47.6 $\rightarrow$ 33.3, Media: 57.1 $\rightarrow$ 28.6).
-This suggests that some web-based tasks require additional step modifications to improve execution accuracy.
-Actor-Critic is the most critical module, as its removal results in the sharpest performance drop across all task groups (Overall SR: 36.2 $\rightarrow$ 13.3).
-We observe that many subtasks are difficult to execute correctly on the first attempt and heavily rely on the Actor-Critic for additional action correction.
-The most drastic declines are seen in Coding (45.5 $\rightarrow$ 0.0), Windows Usage (34.8 $\rightarrow$ 4.3), and Office (44.4 $\rightarrow$ 15.6), indicating that these tasks are more rely on the Actor-Critic module.
-Overall, these results demonstrate that all three modules contribute significantly to the performance, with Actor-Critic being the most impactful.
-Impact of Different MLLM.
-We also investigate the impact of using different MLLMs in our framework.
-As shown in Tab.
-3, we compare three MLLMs across five popular applications.
-Claude-3.5 consistently achieves the highest performance on all five applications, particularly excelling in VSCode and Acrobat tasks, where it attains an SR of over 60%.
-GPT-4o demonstrates stable performance across these applications, while Gemini-2.0 underperforms compared to the other two MLLMs.
-Impact of Instructional Video.
-In Tab.
-4, we study the impact of removing the instructional video by modifying the prompt to include only the user query for generating the initial plan.
-On the three Office applications, we observe a significant performance decline, as these tasks rely more heavily on additional domain knowledge for successful planning.
-In contrast, the MLLM performs relatively well on Win.
-Usage tasks, such as Settings and File Management, where it has more inherent familiarity.
-Conclusion
-In this paper, we take the first step toward comprehensive GUI agent evaluation by introducing a new benchmark, WorldGUI.
-In addition to the standard static testing processes, we incorporate dynamic testing procedures to ensure that WorldGUI effectively captures the complexity and dynamism of real-world GUI environments.
-Furthermore, to enhance GUI automation, we propose a novel agent framework, GUI-Thinker, built upon a critical thinking philosophy and comprising five core components.
-This framework enables the agent to dynamically identify uncommon states and adjust its plans or actions accordingly.
-Finally, we evaluate the latest computer-using agent, Claude-3.5, using our WorldGUI benchmark, demonstrating the effectiveness of GUI-Thinker across a variety of GUI tasks.
-Impact Statement
-This paper introduces WorldGUI and GUI-Thinker, a benchmark and agent framework designed to advance the field of GUI automation.
-Our work has several important societal and research implications.
-Firstly, WorldGUI represents a significant step toward comprehensive GUI agent evaluation by incorporating diverse initial states for each GUI task.
-This approach better reflects real-world conditions, inspiring future GUI-related benchmarks to emphasize realistic GUI testing scenarios in their designs.
-Secondly, we propose a holistic GUI agent framework, GUI-Thinker, which integrates five essential components and an iterative reasoning loop.
-By embedding critical thinking into its module design, GUI-Thinker enhances adaptability in complex and dynamic GUI environments, allowing agents to handle unpredictable interface changes more effectively.
-Our work lays a strong foundation for future research in adaptive, reasoning-driven GUI automation, with potential applications in software testing, accessibility tools, and intelligent user assistance systems.
-Appendix A Comparison with other benchmarks
-In this section, we summarize the data statistics of other benchmarks in Tab.
-5 as the comparison.
-We can observe that the main difference of our WorldGUI is we incorporate various initial states to stimulate the real GUI automation while other benchmarks are not considered.
-Benchmark Softwares Tasks Platform Task Focus Inst. Video GT Plan Various States
-WebArena (Zhou et al., ) 6 812 Web Website Usage
-VisualWebArena (Koh et al., 2024) 3 910 Web Website Usage
-Mobile-Eval (Wang et al., 2024) 10 30 Android OS APP Usage
-APPAgent (Zhang et al., 2023) 10 50 Android OS APP Usage
-OSWorld (Xie et al., 2024) 10 369 Ubuntu+Win. Software Usage
-AssistGUI (Gao et al., 2024) 9 100 Windows Productivity Software Usage
-VideoGUI (Lin et al., 2024a) 10 86 Win. + Web Productivity Software Usage
-WindowAgentArena (Bonatti et al., 2024) 11 154 Windows SoftWare Usage
-WorldGUI 10 315 Win. + Web Productivity Software Usage
-Table 5: Comparison with related benmarks. WorldGUI is a unique benchmark that has the various states for each task to stimulate the real-world agent-computer interactions.
-Why Instructional Video?
-The rationale for using instructional videos is that, in some cases, completing a task requires extensive configuration, making it impractical to provide all necessary information within the input query.
-As illustrated in Fig.
-7, our user query is “Generate a photo of a girl with short brown hair with EasyNegative to improve the quality”, which does not specify how to improve the quality.
-The instructional video will serve as contextual information to guide the agent to complete that.
-Figure 7: An example of using Stable Diffusion for media creation in WorldGUI.
-The left shows the interface of using stable diffusion to create the photo.
-The right shows the user query and GT plan.
-If no instructional video, it is hard to include all the settings
-Appendix B Details of Actor Space
-In this section, we detail the action space used in our WorldGUI.
-Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others.
-Mouse-related actions also specify the target position in the pixel space of the observed screenshot.
-To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI, for controlling mouse and keyboard inputs.
-Each action is represented using the syntax action_type(arguments) as in Tab.
-6.
+Table 2: The action types and their example in WorldGUI.
 Action Type Example
@@
 Key Down and Up keyDown(’shift’)
-Table 6: The action types and its example in WorldGUI.
-Appendix C Data
-(1) Annotators.
+Action Space.
+Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others.
+Mouse-related actions also specify the target position in the pixel space of the observed screenshot.
+To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI [^1] https://pyautogui.readthedocs.io, for controlling mouse and keyboard inputs.
+Each action is represented using the syntax action_type(arguments) as in Table  2.
+2.2 Data Collection
+2.2.1 Data Source
+WorldGUI consists of a broad spectrum of widely-used desktop applications, which can be categorized into five main groups: (i) Office, includes PowerPoint, Word, Excel, and Adobe Acrobat; (ii) Windows Usage, includes System Settings and File Management; (iii) Web Usage, includes the configuration of Youtube and website operations; (iv) Coding, focus on the customization, configuration and editing of Visual Studio Code (VSCode); (v) Media, operating VLC player for video editing and creation.
+2.2.2 Pipeline of Data Construction
+We engaged four annotators and developed the necessary scripts to structure and format the data.
+Additionally, to facilitate ground truth (GT) plan generation and pre-action generation, we implemented simple agent-based methods to collect the relevant data.
+The overall data construction pipeline comprises six steps, as detailed below.
+Raw Video Collection.
+We collect raw videos from the YouTube website as there are a lot of high-quality tutorials for desktop applications with high views.
+For each software, we ask the annotators to watch the videos first and download them via the diversity of software usage.
+Instruction Video Preparation.
+After obtaining the raw videos, we write the script codes to cut the lengthy and noisy videos into the sub-clips (30 seconds to 3 minutes) that serve as the instructional video.
+User Query Generation.
+After obtaining the instructional videos, annotators are asked to manually write user queries corresponding to each video.
+For example, a user query for a task involving File Explorer might be: “Please compress the project.mp4 into an MPEG-4 file optimized in 1080p.”
+Project File Preparation.
+Following the AssistGUI assistgui , we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines OSWorld or Docker environments windowsagentarena .
+This approach guarantees that the testing process begins from a consistent state.
+When combined with pre-actions, it enables augmentation of the same task with various initial states.
+GT Plan Generation.
+We write the script to accept user query $q$ and instructional video $v$ as input and generate the raw plans by agent (powered by GPT-4o).
+Since the raw plans are not flawless, annotators are asked to watch the videos and manually execute the tasks following the raw plans.
+During this process, annotators edit the plans to correct any inaccurate steps or descriptions, ultimately producing the finalized GT plans.
+[…]
+2.3 Evaluation
+WorldGUI employs an execution-oriented evaluation approach followed by AssistGUI assistgui and WindowsAgentArena windowsagentarena by utilizing post-processing scripts to assess task completion.
+Specifically, for tasks like Office work and Web Browsing, we adopt exact matching to compare the differences between the ground-truth (GT) screenshots and the final screenshots.
+For tasks like File Management, which would produce new folders or change the locations of files, etc.
+We create the shell script to check the status of files.
+Table 3: Task category, task activities, and project file of the desktop applications in WorldGUI.
+Category Applications All Task Task Activities Project File Type
+Office PowerPoint 64 Change the content style and layout; Design new effects project.pptx
+Office Word 63 Formatting the content style and layout project.docx
+Office Excel 70 Table formatting; Data management and processing project.xlsx
+Office Adobe Acrobat 66 Automatic add electric signature; Document management project.pdf
+Coding VSCode 56 Code editing; Software configuration vscode.exe
+Windows Usage Settings 69 Advanced personalized and safety settings; ms-Settings
+Windows Usage File Explorer 44 File management: Add, delete, rename, and move files explorer.exe
+Web Usage Web Browser 59 Web operation web browser + URL
+Web Usage Youtube (Online) 61 Video and account configurations web browser + URL
+Media VLC Player 59 Video editing and creation project.mp4
+2.4 Data Statistics
+WorldGUI compiles GUI tasks from 10 widely used applications on the Windows platform, including productivity software such as PowerPoint, Word, Excel, and VSCode.
+A total of 111 meta tasks were collected from these applications, with each task being augmented  5 times based on the task’s functionality, resulting in 500 augmented tasks.
+In total, WorldGUI comprises 611 tasks, and every task has almost 6 variation instances, which is capable of reflecting the real-world interactions of the GUI environment.
+See the details in Table 3 (more details in the Supplementary Material).
+3 WorldGUI-Agent: Thinking before Doing
+Figure 3: WorldGUI-Agent.
+The Planner module receives the user query and an instructional video as input and generates an initial plan.
+This plan is then refined and executed step by step.
+Before each step is passed to the Actor module, it undergoes a Step-Check.
+After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes necessary corrections.
+In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  hong2024cogagent ; cheng2024seeclick ; lin2024showui ; zhang2023appagent ; AgentS .
+The WorldGUI-Agent includes the five fundamental but essential components as in Figure 3 and an Interaction reasoning loop detailed in Algorithm 1.
+We summarize our critical designs in the following:
+$\bullet$ Post-Planning Critique: After the planning phase, a critique module verifies and, if necessary, self-corrects the generated plans to ensure their accuracy.
+$\bullet$ Pre-Action Validation: Before executing each subtask, a validation module determines whether the subtask should be executed.
+This step is crucial, as the current GUI environment may indicate that the subtask is unnecessary or requires modification to align with the current state.
+$\bullet$ Post-Action Evaluation: After each action execution, a mechanism evaluates whether the action was successfully completed before proceeding to the next subtask.
+These critique designs ensure the reliability and adaptability of WorldGUI-Agent in complex GUI environments.
+Figure 4: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic provides necessary corrections.
+3.1 State-Aware Planner
+The State-Aware Planner processes the instructional video $v$ and the user query $q$ generates an initial plan as shown in the left of Figure 4.
+We use the speech recognition model Whisper radford2023robust to translate the video $v$ into the subtitle and then send it to the MLLM for task planning.
+The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task.
+Under each $p_{i}$, there is a list of subtasks $[S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone.
+To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state.
+This additional context allows the agent to output plans that align with the actual state.
+For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again.
+By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs.
+It also avoids the occlusion issue when not seeing the screenshot.
+3.2 Planner-Critic
+Post-Planning Critique.
+The goal of the Planner-Critic is to assess the correctness of the initial plans generated by the State-Aware Planner and provide corrections if needed.
+This module is designed to ensure the accuracy of the plans while leveraging the self-reflection capabilities of MLLMs.
+As illustrated in Figure 4, for each Initial Plan, the output consists of four components:
+(1) <Flag>: Indicates whether the Initial Plan is correct.
+(2) <Feedback>: Identifies the error type, categorized into one of three options: “Wrong Steps,” “Missing Steps,” or “Redundant Steps.”
+(3) <Correction>: Provide the corrected plans if the Flag indicates that the Initial Plan is incorrect.
+(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons.
+As related works, CoT wei2022chain , OpenAI-o1 openaio1 , and Deepseek-R1 deepseekr1 demonstrate that generating reasoning steps along with the answer would enhance the performance.
+3.3 Step-Check
+Pre-Action Validation.
+After the plan assessment, a navigation mechanism is crucial before sending each subtask $S_{t}=S_{j}^{i}$ at the time step $t$ to the Actor module.
+To address this, we designed a new module called Step-Check.
+Through extensive investigation, we discovered that during GUI task testing, perfect execution plans are rarely feasible due to the unpredictable nature of real application environments.
+Most software retains user preferences (e.g., remember the last configuration of user), meaning that when executing a specific task, the plan $p$ generated by the Planner might not align with the actual state of the software.
+Therefore, the model must determine whether to proceed with a subtask $S_{t}$ based on the current state (screenshot: $V_{t}$, metadata: $M_{t}$).
+As illustrated in Figure 5, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into four types:
+(1) <Modify>: Indicates that the subtask should be modified or additional subtasks should be added.
+(2) <Pass>: Indicates that the current subtask is unnecessary and can be skipped.
+Figure 5: Step-Check. This module first checks the step completion status via an MLLM and then navigates current task processing.
+(3) <Continue>: Indicates that the subtask is valid and should be executed as planned.
+(4) <Finished>: Indicates that the subtask has already been completed and requires no further action.
+In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs ##\##Cannot confirm”.
+When this occurs, we design a Region Search module implemented by an LLM.
+This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region.
+It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the original screenshot).
+The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
+3.4 Actor
+The goal of the Actor is to translate natural language subtask $S_{t}$ into executable code $C_{t}$.
+Using an MLLM as the backbone model, the Actor processes metadata $m_{t}$ and screenshot $V_{t}$ as GUI context to generate precise executable actions, such as click(100, 200).
+Additionally, it leverages the history of previous actions as memory to aid in generating subsequent actions.
+The generated actions will be executed in the environment, and then the new screenshot $V_{t+1}$ and metadata $m_{t+1}$ will be captured for the next processing.
+3.5 Actor-Critic
+Figure 6: Actor-Critic. This module includes two parts: task verification and task correction. The design follows the verify-then-correct mechanism.
+[…]
+4 Experiments
+Table 4: Success rate (%) of different agents on WorldGUI.
+Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model.
+Meta represents the meta task, while Aug.
+represents the augmented task.
+Method MLLM Office Win. Usage Web Coding Media Overall
+Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
+Plan-Act Gemini-2.0 8.9 3.2 8.3 3.4 28.6 16.2 18.2 2.2 10.0 2.0 6.9
+Plan-Act GPT-4o 13.3 10.1 8.3 2.3 23.8 11.1 9.1 2.2 10.0 2.0 8.5
+AssistGUI GPT-4o 26.7 16.1 29.2 7.9 33.3 20.2 27.3 11.1 10.0 8.2 16.5
+Computer Use Claude-3.5 28.9 19.3 29.2 14.6 71.4 32.3 54.6 22.2 30.0 6.1 23.6
+WorldGUI-Agent Gemini-2.0 31.1 17.0 20.8 9.0 38.1 29.3 36.4 11.1 20.0 10.2 19.1
+WorldGUI-Agent GPT-4o 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
+WorldGUI-Agent Claude-3.5 57.8 32.6 50.0 19.1 76.2 46.5 54.6 26.7 50.0 18.4 36.0
+Human 88.9 83.5 100.0 89.9 95.2 80.8 81.8 77.8 90.0 85.7 85.3
+Implementation Details.
+We implement the MLLM in our WorldGUI-Agent by using GPT-4o gpt4o (gpt-4o-2024-08-06) by default.
+For the computer mouse and keyboard control, we use the Python library PyAutoGUI.
+Following the AssistGUI assistgui , we use the GUI parser to obtain the position information of elements, e.g., buttons, icons, and text.
+We use some vision foundation models, such as Google OCR, to extract the text.
+By default, we use the center coordinates to represent the location of each element.
+All the testing is under the same screenshot resolution (1920 $\times$ 1080).
+In all experiments, we set the max trials of the Actor-Critic to 3 for light interaction costs.
+For the total trials of each task, we set it to $14\times N+$ 1, where $N$ is set empirically.
+Evaluation.
+Given that our WorldGUI includes 611 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results.
+Metric.
+Following the previous works of OSworld and AssistGUI, we use Success Rate (SR) as the metric.
+Baselines.
+We implement the baseline approach called Plan-Act with different MLLMs as the base model.
+It focuses on investigating the basic capabilities of task planning and action prediction.
+Additionally, we compare our WorldGUI-Agent with two strong approaches: AssistGUI assistgui and Computer Use (Claude 3.5) claude3.5 .
+AssistGUI is a prominent agent framework designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query.
+We implement it by increasing the MLLM to GPT-4o for better performance.
+Computer Use (Claude 3.5) is the leading proprietary model specially designed for autonomous computer use.
+We use the open-source implementation OOTB hu2024dawnguiagentpreliminary as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison.
+We also implement our WorldGUI-Agent with three different MLLMs to illustrate the effectiveness of our proposed universal agent framework.
+4.1 Main Results on WorldGUI
+Table 4 reports the success rates (SR) of different agents and human experts on our WorldGUI benchmark, broken down by task type (Meta vs.
+Aug.) across five categories: Office, Win.
+Usage, Web, Coding, and Media.
+From these results we draw the following main conclusions.
+A large gap remains between agents and humans.
+The best-performing agent (WorldGUI-Agent with Claude-3.5) achieves an overall SR of only 36.0%, which is less than half of the 85.3% attained by human experts.
+This stark contrast underscores the difficulty of our tasks and the need for further advances in desktop GUI automation.
+Agents generalize poorly to augmented tasks.
+Across all methods, performance on Augmentation tasks (which introduce interface or context variations) is substantially lower than on their corresponding Meta tasks.
+For example, Claude-3.5 in the Win.
+Usage category attains 50.0% on Meta tasks but drops to just 19.1% on Aug.
+tasks.
+This highlights the importance of dynamic testing to capture realistic human–computer interaction.
+Desktop applications pose a greater challenge than web tasks.
+Every agent scores higher on Web tasks than on desktop application tasks.
+WorldGUI-Agent with Claude-3.5, for instance, jumps from 76.2% on Web Meta to only 57.8% on Office Meta, and the gap widens on their Augmentation counterparts.
+Thus, desktop GUI automation remain a frontier for computer use research.
+WorldGUI-Agent consistently outperforms a naive Plan-Act baseline.
+By incorporating our three critical modules into the planning and execution loop, WorldGUI-Agent substantially improves success rates over the basic Plan-Act approach.
+Relative to Plan-Act, WorldGUI-Agent raises overall SR by +12.2% with Gemini-2.0, +17.5% with GPT-4o, and +12.4% with Claude-3.5, demonstrating the effectiveness of our design across multiple MLLMs.
+Table 5: Success rate (%) of our WorldGUI-Agent (Full Model) with the ablation of different critical modules.
+Method Office Win. Usage Web Coding Media Overall
+Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
+Full Model 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
+– w/o Planner-Critic 31.1 17.0 20.8 9.0 38.1 25.3 36.4 11.1 20.0 10.2 18.5
+– w/o Step-Check 31.1 19.3 20.8 9.0 33.3 28.3 45.5 13.3 20.0 8.2 19.8
+– w/o Actor-Critic 15.6 7.8 4.2 3.4 28.6 17.2 0.0 8.9 10.0 6.1 9.7
+Table 6: Success rate (%) of our WorldGUI-Agent with the ablation of Instructional Video (Inst. Video).
+Method PPT Word Excel Acrobat VSCode Overall
+Full Model 45.5 36.4 50.0 36.4 45.5 42.9
+w/o Inst. Video 45.5 27.3 25.3 18.2 27.3 28.6
+4.2 Ablation Study
+Impact of different critical modules.
+Table 5 presents the results of an ablation study on the three core components of WorldGUI-Agent across five application categories (Office, Windows Usage, Web, Coding, Media).
+The full model achieves an overall success rate (SR) of 26.0%.
+The effects of removing each component are as follows: Planner-Critic: Eliminating this module reduces overall SR to 18.5% (–7.5%), with substantial drops in Office (42.2% → 31.1%) and Web (47.6% → 38.1%) tasks, indicating its importance for refining initial plans.
+Step-Check: Without step-wise verification, SR decreases to 19.8% (–6.2%).
+The relatively smaller decline on Coding and Win.
+Usage tasks suggest that Step-Check excels at intercepting and correcting multi-step interaction errors.
+Actor-Critic: Removing the action-level critic causes SR to collapse to 9.7% (–16.3%).
+Performance on Coding Meta drops to 0.0% and Windows Usage Meta to 4.2%, highlighting the critical role of reward-driven action correction for action-level GUI operations.
+These results confirm that Planner-Critic, Step-Check, and Actor-Critic each contribute complementary benefits—plan refinement, intermediate validation, and action optimization—that are essential for the robustness and overall effectiveness of WorldGUI-Agent.
+Impact of Instructional Video.
+In Table 6, we study the impact of removing the instructional video by modifying the prompt to include only the user query for generating the initial plan.
+In the Excel applications, we observe a significant performance decline, as their tasks are complex and difficult, and rely more heavily on additional domain knowledge for successful planning.
+In contrast, the MLLM performs relatively well on Win.
+Usage tasks, such as Settings and File Management, are where it has more inherent familiarity.
+These findings underscore the necessity of instructional videos for complex tasks like visual effect design, mirroring how users learning to build a slide often rely on tutorial videos.
+4.3 Results on WindowsAgentArena Benchmark
+Table 7 compares WorldGUI-Agent against four leading agents on the WindowsAgentArena benchmark.
+WorldGUI-Agent achieves a 30.5% overall SR, far surpassing GPT-4V (19.5%), GPT-4o (8.6%), GPT-4o-mini (4.2%), and Phi3-V (3.5%).
+Its gains are most pronounced in desktop categories: Office tasks (4.7% vs.
+0%), Windows System (45.8% vs.
+33.3%), and Windows Utilities (33.3% vs.
+8.3%).
+On web browser, it reaches 53.3%, nearly double GPT-4V’s 27.3%, and on coding tasks, it records 33.3% versus 27.3%.
+In media tasks, WorldGUI-Agent posts 28.6%, closely matching GPT-4V’s 30.3%.
+These results underscore the necessity of integrated planning critique, step-check verification, and action-level feedback.
+These results demonstrate that our framework robustly handles both desktop GUI tasks and dynamic web environments, highlighting its versatility for real-world GUI automation.
+Table 7: Experimental results on WindowsAgentArena windowsagentarena . The reported results are from the windowsagentarena .
+Method Office Web Win. System Coding Media Win. Utils Overall
+Phi3-V 0.0 6.9 8.3 0.0 6.2 0.0 3.5
+GPT-4o-mini 0.0 14.9 8.3 0.0 0.0 0.0 4.2
+GPT-4o 0.0 13.7 29.2 0.0 10.3 0.0 8.6
+GPT-4V 0.0 27.3 33.3 27.3 30.3 8.3 19.5
+WorldGUI-Agent (Ours) 7.0 53.3 45.8 33.3 28.6 33.3 31.2
+Conclusion
+In this paper, we take the first step toward comprehensive GUI agent evaluation by introducing WorldGUI.
+In addition to the standard static testing processes, we incorporate dynamic testing procedures to ensure that WorldGUI effectively captures the complexity and dynamism of real-world GUI environments.
+Furthermore, we propose a universal agent framework, WorldGUI-Agent, built upon the critical thinking principle.
+This framework enables the agent to dynamically identify uncommon states and adjust its plans or actions accordingly.
+Finally, we evaluate WorldGUI-Agent powered by Claude-3.5-Sonnet on WorldGUI and WindowsAgentArena benchmarks, demonstrating the effectiveness across a variety of GUI tasks.
+Limitation and Society Impacts
+In the current implementation of our agent, WorldGUI-Agent, more external tools have not been integrated into the GUI planning and action prediction processes to prioritize computational efficiency.
+Incorporating tools such as web search or file search into the agent’s design could be a valuable future direction to improve performance.
+Additionally, due to the usage of the GUI Parser, which would increase the time costs because of the response speed of experimental desktop computers, it is still a tradeoff between performance and running time in the current GUI domain.
+We consider that if the base MLLM model is specifically trained with stronger planning and grounding ability, the running time would be sped up.
+It is noted that our agent framework is capable of working with any MLLM.
+WorldGUI takes the first step of pushing the GUI automation into the dynamic testing process, as we found that real-world human-computer interactions are dynamic and unpredictable; existing GUI benchmarks fail to capture such dynamics to closely reflect the interactions.
+Our WorldGUI-Agent is a straightforward and universal agent framework by considering incorporates three critical modules to adaptively align the plan and actions with exact environment situations, which would be a good baseline for future agent development.
+For instance, incorporating more tools such as web search or file search into the planning module or action prediction module to realize more challenging tasks.
+Appendix A Related Work
+A.1 GUI Benchmarks
+GUI benchmarks are essential for evaluating the performance and robustness of GUI agents.
+For web applications, WebShop [29], WebArena [34], and WebVoyager [10] focus on creating the GUI tasks in a web browsing scenario.
+In OS environments, OSWorld [28] is a comprehensive benchmark, including various operating systems with real applications.
+Mobile benchmarks such as MobileAgent [23] and AppAgent [31] propose two GUI benchmarks of mobile applications.
+Windows-related benchmarks like AssistGUI [7] and WindowAgentArena [4] propose a list of real tasks in the Windows platform.
+However, these online testing GUI benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments.
+As a result, they are insufficient for comprehensively evaluating GUI agents.
+A.2 GUI Agents
+CogAgent [12] is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick [5] and SeeAct [32] focus on the GUI grounding for enhancing the task performance.
+MobileAgent [23] and AppAgent [31] are proposed to design the agent on the mobile device.
+Ferret-UI [30] is another representative work focusing on enhancing the grounding ability in the IOS platform.
+These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction, but still face limitations in handling dynamic and complicated full GUI tasks.
+Therefore, to enhance GUI automation in dynamic environments, we propose WorldGUI-Agent, which improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes.
+The components comparison of our WorldGUI-Agent and other closely related agents is shown in Table 8.
+Table 8: Comparison with other closely related agents.
+Most existing agents solely focus on post-action evaluation but omit the post-planning critique and pre-action validation in handling dynamic GUI environments.
+Method Post-Planning Critique Pre-Action Validation Post-Action Evaluation
+Mobile-Agent [23]
+Mobile-Agent-V2 [23]
+AssistGUI [7]
+Agent-S [1]
+Mobile-Agent-E [24]
+WorldGUI-Agent (ours)
+A.3 Critical Thinking in Agents
+Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 [19] and Deepseek-R1 [6], have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks.
+The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency.
+One notable LLM-based agent framework, Reflexion [22], demonstrates the effectiveness of self-reflection in solving complex tasks.
+Furthermore, CRITIC [8] integrates external tools into the critique process, leveraging them to improve performance.
+Noticing that the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario.
+Which is not only aims to enhance the reasoning performance but is also indispensable to designing the key module Actor-Critic [15] to ensure task completion.
+A closely related work, AssistGUI [7], integrates a critical module only after the Actor module to evaluate action completion.
+Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor.
+These two modules lead to a universal and fundamental GUI agent framework WorldGUI-Agent which will provide insights for future GUI agent design.
+Appendix B Data
+Annotators.
 In this work, we have four annotators: A, B, C, and D.
@@
 For the 10 desktop applications, we divide the software into four parts, assigning each part to a different annotator.
-For the human tests presented in Tab.
-1, the annotators demonstrate tasks on software that they did not annotate.
-As shown in Tab.
-5, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.
+For the human tests presented in Table 4, the annotators demonstrate tasks on software that they did not annotate.
+As shown in Table 1, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.
+Table 9: The annotation arrangement during the annotation and human testing phases by different annotators.
 Annotators Annotation Phase Human Test Phase
-A PowerPoint, Word, Excel VSCode, Stable Diffusion, Web
-B Adobe Acrobat, Stable Diffusion Excel, Settings
+A PowerPoint, Word, Excel VSCode, VLC Player, Web
+B Adobe Acrobat, VLC Player Excel, Settings
 C Settings, Web PowerPoint, File Explorer, Youtube
 D VSCode, File Explorer Word, Adobe Acrobat
-Table 7: This table shows the software used during the annotation and human testing phases by different annotators.
-(2) Data Construction Pipeline.
-For each task 1) we manually collect the raw videos from the website, 2) then manually cut the raw lengthy videos into sub-clips, 3) ask annotators to manually write the user queries, 4) prepare the project files for each task for reproducibility, 5) generate the ground-truth plans by executing the scripts and agent, 6) conduct the data augmentation by varying the initial state for each task.
-As shown in Fig.
-11, to achieve the above six steps, we invite four annotators and write the necessary scripts to help structure and format the data.
-Additionally, for the GT plan generation and pre-action generation, we build simple agents to obtain the data.
-(3) Project File and Enviroment Configuration.
-In this paper, we utilize project files to ensure the reproducibility of task executions.
-This approach is simpler and more cost-effective compared to using virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024).
-The types of project files for each application are illustrated in Tab.
-8.
-To enable cross-platform usage, such as controlling the computer via a phone or other systems, we implement a frontend and backend system.
-The frontend receives user queries, captures screenshots and metadata, and then sends this information to the backend for API calls and the invocation of local vision foundation models.
-Cross-platform support allows us to deploy autonomous agents in a simple and cost-effective manner.
-(4) Creating Augmented Tasks.
+Figure 7: Pipeline of Data Construction.
+Human: Represents the annotators.
+Code: Refers to the scripts (e.g., Python Code) utilized to achieve the goal.
+Agent: We design an agent built upon the MLLMs to achieve the goal.
+Creating Augmented Tasks.
 In our study, to stimulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks.
@@
 If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
-2) Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings.
+Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings.
 This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps.
-We illustrate one example in Fig.
-8.
+We illustrate one example in Figure 8.
 Here, the meta task and augmented task, have the same user query and instructional video and it will ideally have the same final state.
-We additionally provide more examples about augmenting the meta task in Fig.
-9.
+We additionally provide more examples about augmenting the meta task in Figure 9.
 Figure 8: An example of augmenting one GUI task with manually aug the initial state and then using the execution scripts and corresponding agents to obtain the pre-action for each augmented case.
 Figure 9: We present the examples of conducting the augmentations from the meta task.
-(5) Data Statistic.
-WorldGUI comprises 315 GUI tasks sourced from 10 widely used Windows applications, including PowerPoint, Excel, Word, Adobe Acrobat, VSCode, File Explorer, Settings, Web, Youtube, and Stable Diffusion.
-Here we provide the details about the task activities and project file for each application in Tab.
-8.
-We collected 107 meta tasks from these applications, each of which was augmented 0 to 3 times based on its specific content, resulting in 208 augmented tasks.
-Fig.
-10 presents the task distribution, provides query examples, and illustrates the distribution of meta and augmented tasks across different applications.
-Applications All Task Task Activities Project File Type
-PowerPoint 34 Change the content style and layout; Design new effects. project.pptx
-Word 31 Formatting the content style and layout project.docx
-Excel 35 Table formatting; Data management and processing project.xlsx
-Adobe Acrobat 36 Automatic add electric signature; Document management project.pdf
-VSCode 33 Code editing; Software configuration vscode.exe
-Settings 39 Advanced personalized and safety settings; ms-Settings
-File Explorer 26 File management: Add, delete, rename, and move files. explorer.exe
-Web 32 Web operation web browser + URL
-Youtube 28 Video and account configurations web browser + UR
-Stable Diffusion 21 Media creation with complicated effects Local Run Platform
-Table 8: This table shows all tasks, task activities, and project file of the desktop applications used in WorldGUI.
-Figure 10: Distribution of collect tasks, selected queries, and task amount of WorldGUI.
-We have gathered tasks across 10 desktop applications, focusing on the use of productivity software as well as fundamental computer operations and settings.
-Figure 11: Pipeline of Data Construction.
-Human: Represents the annotators.
-Code: Refers to the scripts (e.g., Python Code) utilized to achieve the goal.
-Agent: We design an agent built upon the MLLMs to achieve the goal.
-Appendix D GUI-Thinker Reasoning Loop Algorithm
+Appendix C Data Statistics
+Augmentation tasks type analysis.
+As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state.
+(2) The human-computer interactions may start from the intermediate state of a specific task.
+We propose to create the augmented tasks for each meta task to stimulate authentic GUI interactions.
+Our augmentations lie in two main groups: (1) stimulating the intermediate states and (2) introducing diverse initial states.
+We divide the two groups into three exact types: Add-step, Trim-step, and Adjust-step.
+For Add-step, it represents various unrelated state augmentations to stimulate the scenario that we may start the agent-computer interactions in another unrelated task or interfaces, the agent should replan the task to add necessary steps.
+For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state.
+For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu.
+Most of the time, it would not require new steps to return to the target task progress.
+This augmentation may mislead the agent in state understanding, making them jump or miss the key steps.
+As shown in Figure 11, the manually created augmentations mainly belong to the add-step.
+Adjust-step could be the second-largest application, except for the File Explorer.
+Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
+Task difficulty analysis.
+Figure 12 shows the distribution of the task difficulty across desktop applications.
+We annotated the task difficulty level based the subjective software usage experience.
+The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging.
+The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels.
+By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents.
+According to VSCode and File Explorer, and YouTube applications, the tasks are easier than in other applications.
+Overall, the task difficulty of created data is diverse across different applications, and there is still a need for stronger agents focusing on handling desktop-oriented GUI tasks.
+The task details about the user query and the pre-actions are included in the metadata JSON file in the supplementary materials.
+The project file, instruction video, and augmentation files can be found in the provided data link.
+Figure 10: Distribution of Software taxonomy and the distribution of task length.
+Figure 11: The distribution of different augmentation types.
+Figure 12: The distribution of different task difficulty.
+Table 10: Task statistics of WorldGUI.
+Category Software Meta Count Aug Count Meta-Task Length (Avg.)
+Office Adobe Acrobat 11 55 7.1
+Office Excel 12 58 4.1
+Office PowerPoint 11 53 5.1
+Office Word 11 52 5.1
+Coding VSCode 11 45 4.4
+Windows Usage File Explorer 10 34 7.1
+Windows Usage Settings 14 55 5.1
+Web Usage Web Browser 10 49 7.8
+Web Usage YouTube 11 50 4.8
+Media VLC Player 10 49 10.8
+Total (Average) 111 500 6.0
+Appendix D Detailed Experimental Results
+Table 11 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena [4] benchmark.
+The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks.
+As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
+Table 11: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena [4].
+Domain Application #Tasks #Successes SR (%)
+Web Browsing chrome 17 11 64.71
+Windows Utilities clock 4 2 50.00
+Windows System file_explorer 19 7 36.84
+Office libreoffice_calc 24 1 4.17
+Office libreoffice_writer 19 2 10.53
+Windows Utilities microsoft_paint 3 1 33.33
+Web Browsing msedge 13 5 38.46
+Windows Utilities notepad 2 1 50.00
+Windows System settings 5 4 80.00
+Media & Video vlc 21 6 28.57
+Coding vs_code 24 8 33.33
+Windows Utilities windows_calc 3 0 0.00
+Overall 154 48 31.17
+Appendix E Computational Costs Discussion
+The average number of execution steps and tokens consumed are shown below Table 12.
+The execution steps are calculated based on our experimental log files, while the token costs are sampled from representative tasks in each category by taking Actor module as an example.
+Table 12: Average execution steps and token costs on different software.
+Application category Average execution steps Input tokens per step (Actor) per task Output tokens per step (Actor) per task
+Office $\sim$23 2350 212
+Win. Usage $\sim$20 1929 108
+Web $\sim$17 1637 84
+Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a Laptop with AMD Ryzen 7 4800H CPU in Table 13.
+Task length: 6 (Generated by Planner+Planner-Critic).
+Total steps: 18, total time: 335.84s, total agent execution time: 412.37s.
+Since desktop GUI is in its early stages, computational costs are currently unavoidable.
+Even OpenAI’s Deep Research takes over 10 minutes in daily use.
+According to the post from OpenAI’s Operator, achieving 38.1% on OS-World requires over 100 steps, which is costly as well.
+In summary, it is still a tradeoff between performance and time costs in GUI automation.
+Table 13: Time costs of each module.
+Subtask Index Executed Modules Time (seconds)
+0 Planner+Planner-Critic 12.53
+1 Step-Check+Actor 21.23
+1 Actor-Critic 13.07
+2 Step-Check+Actor 17.81
+2 Actor-Critic 11.86
+3 Step-Check+Actor 18.43
+3 Actor-Critic 11.72
+4 Step-Check+Actor 22.49
+4 Actor-Critic 25.31
+4 Actor-Critic 10.82
+5 Step-Check+Actor 17.01
+5 Actor-Critic 22.36
+5 Actor-Critic 26.85
+5 Actor-Critic 18.99
+6 Step-Check+Actor 21.79
+6 Actor-Critic 23.37
+6 Actor-Critic 21.12
+6 Actor-Critic 19.08
+Appendix F Examples of Augmentations
+In this section, we present several augmentation examples in Figures 13, 14, 15, 16, 17, 18, 19.
+It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step.
+For instance, in Figure 13, our augmentation is about click on Data tab in the ribbon, in the default software state, the Merge & Center button exhibit in the Home tab, there is no need to click on Home tab, after our augmentations, the agent should add a new task “Click on Home Tab” before it click on the Merge & Center button.
+Similarly, in Figure 18, the text editing buttons are under the Home Tab, if we augment the initial state with other Tab like Animation Tab, after the first step “Select the text ’US SUBMARINE DAY’ ”, the agent should add a new step like “Click on Home Tab” back to the default state for task execution.
+Except for adding new steps, we also present an example about adjust the step in Figure 14, as the target is about merging cells A1 to K1, we augment the initial state by selecting A2 to K2.
+Such a slight difference may mislead the agent to perceive such a minor difference, and the agent may jump the first step about selecting the correct cells lead to finally unsucess.
+In Figure 15 and Figure 19, we show two examples of introducing pop-up window in the initial state which require the agents accurately identify the pop-up windows and correctly close it by replanning the task based on the visual screenshot not only strictly planning based on inherited knowledge or the instructinal videos.
+In Figure 16, we show an example of changing the interface by clicking the Data tab to hide the Merge & Center button under the Home tab.
+In Figure 17, we complete the first step about selecting A1 to K1, which requires the agent to jump this step to reduce the time costs.
+Figure 13: Augmented example of an Excel Task.
+Figure 14: Augmented example of an Excel Task.
+Figure 15: Augmented example of an Excel Task.
+Figure 16: Augmented example of an Excel Task.
+Figure 17: Augmented example of an Excel Task.
+Figure 18: Augmented example of a PowerPoint Task.
+Figure 19: Augmented example of a PowerPoint Task.
+Appendix G WorldGUI-Agent Reasoning Loop Algorithm
 In this section, we provide the details of our reasoning loop algorithm in Algorithm 1.
-Algorithm 1 GUI-Thinker Reasoning Loop Algorithm
+Algorithm 1 WorldGUI-Agent Reasoning Loop Algorithm
   Input: State $s$, Action Code $C$, Screenshot $V$, Metadata $m$, Current subtask $S$, Critic_count $z$
@@
   end while
-Appendix E Qualitative Results
-(1) In Fig.
-12, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code.
+Appendix H Qualitative Results
+(1) In Figure 20, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code.
 Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
-Figure 12: We show one successful prediction of our GUI-Thinker.
+Figure 20: We show one successful prediction of our WorldGUI-Agent.
 Figure 21: We show two examples of using GUI Parser to obtain the element position information.
-(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Fig.
-14, Fig.
-15, and Fig.
-16.
+(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 22, Figure 23, and Figure 24.
 These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
@@
 1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application.
-As shown in the left of Fig.
-17, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden.
+As shown in the left of Figure 25, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden.
 Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area.
-2) As shown in the right of Fig.
-17, the model has difficulty dragging a bar to achieve the desired value.
-3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Fig.
-18.
+2) As shown in the right of Figure 25, the model has difficulty dragging a bar to achieve the desired value.
+3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Figure 26.
 The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot.

### v4

@@
 Abstract
-GUI agents have achieved outstanding performance in GUI element grounding.
-However, planning remains highly challenging, especially due to the sensitivity to the initial state of the environment.
-Specifically, slight differences in the initial state-such as the target software not being open or the interface not being in its default state, often lead to planning errors.
-This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it.
-To address this gap, we introduce WorldGUI, a comprehensive GUI benchmark containing tasks across ten widely used desktop and web applications (e.g., PowerPoint, VSCode, Acrobat), each instantiated with diverse initial states to simulate authentic human–computer interactions.
-Complementing this, we propose WorldGUI-Agent, a universal framework that unifies three core modules: Planner-Critic for high-level plan refinement, Step-Check for intermediate verification, and Actor-Critic for action-level optimization to proactively detect and correct errors.
-Experimental evaluation shows that WorldGUI-Agent outperforms the outstanding existing model (Claude-3.5 Computer Use) by 12.4% in success rate on WorldGUI, and achieves a 31.2% overall success rate on WindowsAgentArena, surpassing the prior state-of-the-art by 11.7%.
-Our analysis further reveals that dynamic augmentation tasks and desktop environments pose substantial hurdles, underscoring the necessity of adaptive planning and feedback-driven execution for advancing real-world GUI automation.
+Recent progress in GUI agents has substantially improved visual grounding, yet robust planning remains challenging, particularly when the environment deviates from a canonical initial state.
+In real applications, users often invoke assistance mid-workflow, where software may be partially configured, steps may have been executed in different orders, or the interface may differ from its default setup.
+Such task-state variability is pervasive but insufficiently evaluated in existing GUI benchmarks.
+To address this gap, we introduce WorldGUI, a benchmark covering ten widely used desktop and web applications with tasks instantiated under diverse, systematically constructed initial states.
+These variations capture realistic human–computer interaction settings and enable diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle non-default contexts.
+We further present WorldGUI-Agent, a simple and model-agnostic framework that organizes planning and execution around three critique stages, improving reliability in dynamic environments.
+Experiments demonstrate that state-of-the-art GUI agents exhibit substantial performance degradation under non-default initial conditions, revealing limited robustness and fragile planning behaviors.
+Our benchmark and framework provide a foundation for developing more adaptable and reliable GUI agents.
 The code and data are available at https://github.com/showlab/WorldGUI.
-00footnotetext: Corresponding author.
-Figure 1: Software taxonomy of WorldGUI and the performance comparison of GUI agents.
-The left shows 5 main groups and 10 software in our WorldGUI.
-The right shows that WorldGUI-Agent surpasses previous SOTA GUI agents significantly.
 1 Introduction
-Graphical User Interface (GUI) automation has become a prominent research area, driven by the need to enhance user productivity.
-This domain encompasses software usage, file management, office design, coding, and web browsing.
-Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o gpt4o and Claude-3.5 claude3.5 , GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to enhance productivity efficiency.
-GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition he2016deep and visual question answering antol2015vqa ; VQAv2 .
-However, current online GUI benchmarks such as WebArena webarena , WebVoyager webvoyager , and WindowsAgentArena windowsagentarena do not capture this dynamism.
-Currently, most GUI benchmarks OSWorld ; windowsagentarena ; assistgui ; webarena ; visualwebarena ; webvoyager focus on initial and final states, measuring success rates but overlooking the state variety in real GUI scenarios.
-These benchmarks often ignore situations where: (1) The software interface is not in its default state.
-(2) The human-computer interactions may start from the intermediate state of a specific task.
-(3) Differences in agent robustness, where agents with the same low success rate (e.g., 20%) may vary in their ability to self-reflection, but these abilities cannot be measured in a static setting.
-As a result, these benchmarks fail to comprehensively assess the GUI agents.
-Figure 2: WorldGUI.
-The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions.
-The pre-actions lead to different initial states.
-The key characteristic of our WorldGUI is the various initial states of the same task to stimulate the real-world testing process.
-The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
-In this paper, we take the first step toward comprehensive GUI evaluation by designing GUI tasks with various initial states.
-We consider that the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default initial conditions, allowing tasks to start from intermediate states where users may seek assistance at any point.
-(2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing plans or introducing new steps to ensure task execution.
-By incorporating these situations into the benchmark design, WorldGUI better mirrors real-world GUI interactions, enabling a more accurate and thorough assessment of GUI agent capabilities.
-Specifically, WorldGUI embraces 10 widely-used desktop applications with 611 tasks in total.
-For each task, we create a user query, an instructional video, and the corresponding project file.
-We engaged four trained annotators skilled in using these applications for annotation.
-To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions.
-In addition, we introduce a new GUI agent framework, WorldGUI-Agent, which builds upon critical thinking design principle, an aspect less emphasized in previous GUI agents hong2024cogagent ; cheng2024seeclick ; autowebglm ; AgentS ; osatlas .
-In dynamic GUI environments, application settings may not be in default configurations.
-This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy.
-Through our analysis of real-world GUI scenarios, we identify three design principles for GUI agents: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation.
-We argue that these components are fundamental and universal for GUI agents.
-To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in the online GUI testing and propose a new benchmark called WorldGUI; (2) We introduce WorldGUI-Agent, a fundamental and universal GUI framework that incorporates critical thinking into the overall agent design, providing valuable insight and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling GUI tasks (see Figure 1).
-2 WorldGUI Benchmark
-Table 1: Comparison with other interactive GUI benchmarks.
+Table 1: Comparison with other online GUI benchmarks.
 WorldGUI is a unique benchmark that embraces diverse initial states and better reflects the authentic interactions in GUI scenarios.
 Env?: Indicates whether an environment is required to be deployed.
-Benchmarks Softwares Tasks Platform Env? Inst. Video? GT Plan Diverse Contextual
-Init. State? Variability?
-WebArena webarena 6 812 Web Yes
-VisualWebArena visualwebarena 3 910 Web Yes
-WebVoyager webvoyager 15 643 Web Yes
-AutoDroid AutoDroid 13 158 Android OS Yes
-AndroidWorld androidworld 20 116 Android OS Yes
-AgentStudio zheng2025agentstudio 9 205 Desktop + Web Yes
-Mobile-Eval wang2024mobile 10 30 Android OS Yes
-APPAgent zhang2023appagent 10 50 Android OS Yes
-OSWorld OSWorld 10 369 Desktop Yes
-AssistGUI assistgui 9 100 Windows No
-WindowAgentArena windowsagentarena 11 154 Windows Yes
+Benchmark Software Task Platform Env? Inst. Video? GT Plan Task State
+variability?
+WebArena (Zhou et al., ) 6 812 Web Yes
+VisualWebArena (Koh et al., 2024) 3 910 Web Yes
+WebVoyager (He et al., 2024) 15 643 Web Yes
+AutoDroid (Wen et al., 2024) 13 158 Android OS Yes
+AndroidWorld (Rawles et al., 2024) 20 116 Android OS Yes
+AgentStudio (Zheng et al., 2025) 9 205 Desktop + Web Yes
+Mobile-Eval (Wang et al., 2024) 10 30 Android OS Yes
+APPAgent (Zhang et al., 2023) 10 50 Android OS Yes
+OSWorld (Xie et al., 2024) 10 369 Desktop Yes
+AssistGUI (Gao et al., 2024) 9 100 Windows No
+WindowAgentArena (Bonatti et al., 2024) 11 154 Windows Yes
 WorldGUI 10 611 Win. + Web No
-2.1 Task Formulation
-GUI Automation Definition.
-The GUI automation task can be considered a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$ with state space $\mathcal{S}$, observation $\mathcal{O}$, action space $\mathcal{A}$, transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, and reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
-In our setting, given a natural language query $q$, eg., Format the slide background with gradient fill that describes a specific task in high-level, along with an instructional video $v$ as a supplement that more detailed illustrates how to complete it, the agent first get the observation $o_{t}\in\mathcal{O}$ from the state $s_{t}\in\mathcal{S}$ in the execution environment and then generate the executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in\mathcal{S}$ and a new observation $o_{t+1}\in\mathcal{O}$.
-The process repeats until the task is finished or fails.
-The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step ,indicating the task completion status.
-WorldGUI Task Definition.
-As illustrated in Figure  2, to achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case.
-This is accomplished through the use of pre-actions, which consist of a sequence of executable code to initialize tasks from different initial states.
-With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios.
-We additionally summarize the differences between WorldGUI and other close interactive benchmarks in Table 1.
-Observation Space.
-The observation space $\mathcal{O}$ indicates the information of the operating system (OS) available to the agent in each state $s_{t}$.
-In this paper, we follow the previous work of AssistGUI assistgui , encompassing two types of information: metadata $m_{t}$ from the application and screenshot $V_{t}$ of the current state $s_{t}$.
-The metadata mainly includes the layout of panels and UI trees.
-The screenshot $V_{t}$ offers holistic visual information of the current state used for planning and action generation.
-Table 2: The action types and their example in WorldGUI.
-Action Type Example
-Mouse Movement moveTo(120, 200)
-Mouse Clicks click(200, 300)
-Keyboard Type write(’classes’)
-Hotkey hotkey(’ctrl’, ’a’)
-Scrolling scroll(-100)
-Drag dragTo(120, 220, 2)
-Mouse Down and Up mouseDown(); mouseUp()
-Press Keys press(’delete’)
-Key Down and Up keyDown(’shift’)
-Action Space.
-Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others.
-Mouse-related actions also specify the target position in the pixel space of the observed screenshot.
-To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI [^1] https://pyautogui.readthedocs.io, for controlling mouse and keyboard inputs.
-Each action is represented using the syntax action_type(arguments) as in Table  2.
-2.2 Data Collection
-2.2.1 Data Source
+2 Introduction
+Graphical User Interface (GUI) automation has emerged as a promising direction for productivity-oriented AI systems.
+With the rapid advancement of Multimodal Large Language Models (MLLMs) such as GPT-5 (Singh et al., 2025) and Claude-4-Sonnet (Anthropic, 2025), GUI agents have demonstrated the potential to assist users in software usage, file management, office design, coding, and web browsing, reducing repetitive work and improving workflow efficiency.
+Unlike traditional computer vision tasks such as image recognition (He et al., 2016) or visual question answering (Antol et al., 2015; Goyal et al., 2017), GUI automation operates in a highly dynamic environment where the interface state evolves as a result of user actions, system configurations, or partial task progress.
+However, existing online GUI benchmarks, including WebArena (Zhou et al., ), WebVoyager (He et al., 2024), OSWorld (Xie et al., 2024), AssistGUI (Gao et al., 2024), and WindowsAgentArena (Bonatti et al., 2024), primarily evaluate agents from a single, canonical initial state and measure success only at the end of the trajectory.
+Such static evaluation overlooks several common situations in real-world usage: (1) software interfaces are often not in their default configurations, (2) users frequently request assistance from intermediate states of partially completed tasks, and (3) agents with the same low success rate (e.g., 20%) may differ substantially in their robustness, planning stability, and self-correction behavior differences that cannot be revealed under a single start-state protocol.
+In parallel, robustness-oriented GUI benchmarks such as GUI-Robust (Yang et al., 2025) examine how agents respond to interface-level perturbations, including layout shifts, appearance variations, and transient pop-ups.
+These benchmarks vary the visual presentation while keeping the task goal and initial state fixed, providing valuable insights into perception robustness.
+In contrast, WorldGUI does not alter UI skin, resolution, or layout; instead, it targets task-progress–level variation by modifying the executed steps themselves (Add-, Trim-, and Adjust-step).
+This form of task-state variability, such as missing, redundant, or reordered steps, poses challenges fundamentally different from interface perturbations, making the two robustness dimensions complementary and largely orthogonal.
+To complement existing benchmarks, we focus on robustness to intermediate starting states, a ubiquitous but under-explored aspect of GUI interactions.
+We introduce WorldGUI, a benchmark that systematically constructs diverse initial states for each task by applying controlled pre-actions.
+These pre-actions modify task progress itself (e.g., steps partially completed, over-executed, or executed in a different order), allowing tasks with the same goal to begin from multiple realistic contexts.
+This design mirrors how real users summon assistance in the middle of workflows and enables diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle contextual variability.
+WorldGUI covers 10 widely used desktop applications and provides 611 tasks, each paired with a user query, instructional video, and project file.
+To construct meaningful state variations, trained annotators first demonstrate ground-truth (GT) plans, after which we generate augmented initial states via pre-actions.
+In addition, we propose WorldGUI-Agent, a simple and general GUI agent framework grounded in critical-thinking principles, an aspect that has received relatively limited attention in prior GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2025a; Wu et al., 2024).
+Dynamic GUI environments often deviate from expected states due to user behavior or system configurations, making it essential for agents to detect mismatches and adapt their plans.
+Through an analysis of real-world GUI scenarios, we distill three design principles that we argue are fundamental for GUI automation: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation.
+These components enable WorldGUI-Agent to refine plans, verify step prerequisites, and evaluate action outcomes, thereby improving reliability in dynamic settings.
+Experiments reveal three key observations: (1) a substantial gap remains between humans and current agents, with models such as UI-Tars and GPT-5.1 still struggling on our curated desktop GUI tasks; (2) performance degrades sharply under non-default initial states, highlighting limited robustness and fragile planning capabilities; and (3) WorldGUI-Agent significantly improves robustness and overall task success across WorldGUI and WindowsAgentArena, even when built upon weaker base models and without prompt tuning, demonstrating the importance of multi-level critique for dynamic GUI automation (see Section 6).
+3 Related Work
+Figure 1: WorldGUI.
+Left: WorldGUI creates pre-actions for each meta task, leading to different initial states.
+It successfully reflects the real-world human-computer interaction process.
+Right: components in WorldGUI.
+3.1 GUI Benchmarks
+GUI benchmarks provide the foundation for evaluating computer-use agents across web, desktop, and mobile environments.
+Early interactive benchmarks such as WebShop (Yao et al., 2022), WebArena (Zhou et al., ), and WebVoyager (He et al., 2024) center on browser tasks with long-horizon goals.
+OSWorld (Xie et al., 2024) and mobile benchmarks, including MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) broaden coverage to multi-application desktop and mobile ecosystems.
+AssistGUI (Gao et al., 2024) and WindowsAgentArena (Bonatti et al., 2024) provide online execution environments for Windows-based tasks.
+Recent benchmarks emphasize scalability and online evaluation.
+UIExplorer (Nica et al., 2025) offers a large-scale environment for GUI exploration in both structured-DOM and pure screen modes.
+GUI-World (Chen et al., 2025) provides multi-application desktop tasks with online execution and standardized initial states.
+Robustness-oriented suites such as GUI-Robust (Yang et al., 2025) examine agents under interface-level perturbations (e.g., layout shifts, visual noise, pop-ups), focusing on perception and action robustness.
+While these benchmarks significantly advance GUI automation, they overwhelmingly assume a single canonical initial state for each task.
+They do not explicitly evaluate task-state variability, such as partially completed workflows, missing or redundant steps, or alternative intermediate contexts common in real user workflows.
+Our benchmark, WorldGUI, complements existing online environments by systematically constructing diverse initial states for the same task goal via controlled pre-actions, enabling diagnostic evaluation of robustness, recoverability, and planning stability in dynamic GUI environments.
+3.2 GUI Agents
+Recent progress in GUI agents spans perception, grounding, planning, and full computer-use control.
+CogAgent (Hong et al., 2024) improves GUI perception, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on grounding UI elements for more reliable action prediction.
+MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) extend these capabilities to mobile environments.
+On desktop and OS-level tasks, AssistGUI (Gao et al., 2024) adopts a Plan-Act pipeline with a single post-action checker.
+Agent-S (Agashe et al., 2025a) introduces experience-augmented hierarchical planning with a Manager-Worker architecture, a self-evaluator, and an Agent-Computer Interface (ACI).
+Agent-S2 (Agashe et al., 2025b) further incorporates specialist grounding (Mixture-of-Grounding) and Proactive Hierarchical Planning, achieving strong performance across OSWorld, WindowsAgentArena, and AndroidWorld.
+Beyond inference-time frameworks, several works train GUI agents or their components.
+Aguvis (Xu et al., 2024) and UGround (Gou et al., 2025) train pure-vision or universal grounding models at scale.
+AgentTrek (Xu et al., 2025) synthesizes multi-step trajectories from web tutorials for supervised training, and SE-GUI (Yuan et al., 2025) applies reinforcement learning to strengthen grounding robustness.
+While these approaches enhance perception, grounding, or hierarchical planning, they typically employ only a single verification stage (e.g., self-evaluation or post-action checking).
+In contrast, our framework WorldGUI-Agent focuses on structuring the verification process itself for dynamic GUI environments.
+WorldGUI-Agent introduces three complementary critique stages: Post-Planning Critique, Pre-Action Validation, and Post-Action Evaluation, enabling agents to detect state mismatches, assess step readiness, and repair execution failures.
+This multi-level critique design is orthogonal to hierarchical planners and training-based pipelines, and can be combined with them without additional model training.
+A component-wise comparison is provided in Table 3.
+4 WorldGUI Benchmark
+Following prior work (Xie et al., 2024; Zhou et al., ), we model GUI automation as a partially observable Markov decision process (POMDP).
+A complete specification of the POMDP components and the execution environment is provided in Appendix A.
+4.1 Data Source
 WorldGUI consists of a broad spectrum of widely-used desktop applications, which can be categorized into five main groups: (i) Office, includes PowerPoint, Word, Excel, and Adobe Acrobat; (ii) Windows Usage, includes System Settings and File Management; (iii) Web Usage, includes the configuration of Youtube and website operations; (iv) Coding, focus on the customization, configuration and editing of Visual Studio Code (VSCode); (v) Media, operating VLC player for video editing and creation.
-2.2.2 Pipeline of Data Construction
+Table 2: Task category, task activities, and project file of the desktop applications in WorldGUI.
+Category Applications All Task Avg. Length Task Activities Project File Type
+Office PowerPoint 64 7.1 Change the content style and layout; Design new effects project.pptx
+Office Word 63 4.1 Formatting the content style and layout project.docx
+Office Excel 70 5.1 Table formatting; Data management and processing project.xlsx
+Office Adobe Acrobat 66 5.1 Automatic add electric signature; Document management project.pdf
+Coding VSCode 56 4.4 Code editing; Software configuration vscode.exe
+Windows Usage Settings 69 7.1 Advanced personalized and safety settings; ms-Settings
+Windows Usage File Explorer 44 5.1 File management: Add, delete, rename, and move files explorer.exe
+Web Usage Web Browser 59 7.8 Web operation web browser + URL
+Web Usage Youtube (Online) 61 4.8 Video and account configurations web browser + URL
+Media VLC Player 59 10.8 Video editing and creation project.mp4
+Total (Average) 611 6.0
+Figure 2: The left shows software taxonomy in WorldGUI. The right shows the distribution of task length.
+4.2 Pipeline of Data Construction
 We engaged four annotators and developed the necessary scripts to structure and format the data.
+To reduce annotator-style variance, all annotators followed a standardized annotation protocol.
 Additionally, to facilitate ground truth (GT) plan generation and pre-action generation, we implemented simple agent-based methods to collect the relevant data.
@@
 Project File Preparation.
-Following the AssistGUI assistgui , we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines OSWorld or Docker environments windowsagentarena .
+Following the AssistGUI (Gao et al., 2024), we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines (Xie et al., 2024) or Docker environments (Bonatti et al., 2024).
 This approach guarantees that the testing process begins from a consistent state.
@@
 GT Plan Generation.
-We write the script to accept user query $q$ and instructional video $v$ as input and generate the raw plans by agent (powered by GPT-4o).
-Since the raw plans are not flawless, annotators are asked to watch the videos and manually execute the tasks following the raw plans.
-During this process, annotators edit the plans to correct any inaccurate steps or descriptions, ultimately producing the finalized GT plans.
-[…]
-2.3 Evaluation
-WorldGUI employs an execution-oriented evaluation approach followed by AssistGUI assistgui and WindowsAgentArena windowsagentarena by utilizing post-processing scripts to assess task completion.
-Specifically, for tasks like Office work and Web Browsing, we adopt exact matching to compare the differences between the ground-truth (GT) screenshots and the final screenshots.
-For tasks like File Management, which would produce new folders or change the locations of files, etc.
-We create the shell script to check the status of files.
-Table 3: Task category, task activities, and project file of the desktop applications in WorldGUI.
-Category Applications All Task Task Activities Project File Type
-Office PowerPoint 64 Change the content style and layout; Design new effects project.pptx
-Office Word 63 Formatting the content style and layout project.docx
-Office Excel 70 Table formatting; Data management and processing project.xlsx
-Office Adobe Acrobat 66 Automatic add electric signature; Document management project.pdf
-Coding VSCode 56 Code editing; Software configuration vscode.exe
-Windows Usage Settings 69 Advanced personalized and safety settings; ms-Settings
-Windows Usage File Explorer 44 File management: Add, delete, rename, and move files explorer.exe
-Web Usage Web Browser 59 Web operation web browser + URL
-Web Usage Youtube (Online) 61 Video and account configurations web browser + URL
-Media VLC Player 59 Video editing and creation project.mp4
+Given a user query $q$ and instructional video $v$, we automatically generate an initial draft plan using a GPT-4o–based agent.
+Annotators then watch the video and execute the task while reviewing the draft plan, correcting any inaccurate or missing steps to obtain the final GT plan.
+GT plans are used solely for constructing pre-actions and are never provided to or used by any evaluated agent.
+Pre-Actions Generation.
+To vary the task, we propose introducing pre-actions before the task begins.
+These pre-actions are created by annotators and involve corresponding scripts and agents.
+They are written in Python code, for example: from pyautogui import click, $\n rightClick$(800,400).
+The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state.
+This approach addresses scenarios where users may invoke GUI assistant at any time.
+For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu.
+If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
+2) Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings.
+This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding necessary new steps.
+See example in Figure 5.
+Each pre-action was manually verified to ensure the task remained solvable.
 4.3 Data Statistics
-WorldGUI compiles GUI tasks from 10 widely used applications on the Windows platform, including productivity software such as PowerPoint, Word, Excel, and VSCode.
+Overall.
+We present the detailed task taxonomy and lengths about WorldGUI in Figure 2 and Table 2.
 A total of 111 meta tasks were collected from these applications, with each task being augmented  5 times based on the task’s functionality, resulting in 500 augmented tasks.
 In total, WorldGUI comprises 611 tasks, and every task has almost 6 variation instances, which is capable of reflecting the real-world interactions of the GUI environment.
-See the details in Table 3 (more details in the Supplementary Material).
-3 WorldGUI-Agent: Thinking before Doing
+All data are available for reference.
+[^1] https://huggingface.co/datasets/anonymousABC/WorldGUI-Bench
+Figure 3: The distribution of different augmentation types.
+Figure 4: The distribution of different task difficulty.
+Figure 5: An example of augmenting one GUI task with manually aug the initial state and then using the execution scripts and corresponding agents to obtain the pre-action for each augmented case.
+Augmentation tasks type analysis.
+As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state.
+(2) The human-computer interactions may start from the intermediate state of a specific task.
+Our augmentations lie in two main groups: (1) simulating the intermediate states and (2) introducing diverse initial states.
+We divide the two groups into three types: Add-step, Trim-step, and Adjust-step.
+For Add-step, it represents various unrelated state augmentations to simulate the scenario that we may start the task in another unrelated task or interfaces, the agent should replan the task to add necessary steps.
+For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state.
+For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu.
+Most of the time, it would not require new steps to return to the target task procedure.
+This augmentation may mislead the agent in state understanding, making them jump or miss the key steps.
+As shown in Figure 3, the manually created augmentations mainly belong to the add-step.
+Adjust-step could be the second-largest part.
+Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
+Task difficulty analysis.
+Figure 4 shows the distribution of the task difficulty across desktop applications.
+We annotated the task difficulty level based the subjective software usage experience.
+The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging.
+The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels.
+By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents.
+Overall, the task difficulty of our created data is diverse across different applications.
+5 WorldGUI-Agent
+Table 3: Comparison with closely related GUI agents.
+Method Post-Plan Pre-Action Post-Action
+Mobile-Agent / V2
+AssistGUI
+Agent-S / S2
+Mobile-Agent-E
+WorldGUI-Agent (ours)
+To study robustness under the dynamically perturbed initial states of WorldGUI, we adopt a lightweight, non-training agent framework.
+As summarized in Table 3, most existing agents rely on a single post-action evaluation module.
+Rather than proposing a new algorithm, our aim is simply to structure the standard perception–planning–execution loop into clearer verification stages.WorldGUI-Agent introduces three small modules at different abstraction levels: Planner-Critic refines the draft plan; Step-Check uses the screenshot and metadata to decide whether a plan step should be executed, skipped, or rewritten; Actor-Critic validates each action using the current observation and recent action history.
+Compared to prior self-reflection approaches (e.g., Reflexion, CRITIC), which apply a single global critique, our checks are lightweight and distributed across plan, step, and action levels to better handle task-progress variation.
+Full module specifications and I/O formats are included in Appendix B.
+6 Experimental Results
+Table 4: Success rate (%) of different agents on WorldGUI.
+Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model.
+Meta represents the meta task, while Aug.
+represents the augmented task.
+Method Office Win. Usage Web Coding Media Overall
+Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
+Plan-Act w/ Gemini-2.0 8.9 3.2 8.3 3.4 28.6 16.2 18.2 2.2 10.0 2.0 6.9
+Plan-Act w/ GPT-4o 13.3 10.1 8.3 2.3 23.8 11.1 9.1 2.2 10.0 2.0 8.5
+AssistGUI w/ GPT-4o 26.7 16.1 29.2 7.9 33.3 20.2 27.3 11.1 10.0 8.2 16.5
+CCU w/ Claude-3.5-Sonnet 28.9 19.3 29.2 14.6 71.4 32.3 54.6 22.2 30.0 6.1 23.6
+UI-TARS-1.5 28.9 16.1 12.5 2.2 28.6 9.1 36.7 6.7 0.0 0.0 12.3
+Agent S2 33.3 16.5 70.8 59.6 52.4 45.5 45.5 37.8 20.0 16.3 34.2
+WorldGUI-Agent (Ours)
+w / Gemini-2.0 31.1 17.0 20.8 9.0 38.1 29.3 36.4 11.1 20.0 10.2 19.1
+w / GPT-4o 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
+w / Claude-3.5-Sonnet 57.8 32.6 50.0 19.1 76.2 46.5 54.6 26.7 50.0 18.4 36.0
+w / GPT-5.1 64.4 37.6 62.5 40.4 85.7 53.5 63.6 46.6 60.0 26.5 45.8
+Human 88.9 83.5 100.0 89.9 95.2 80.8 81.8 77.8 90.0 85.7 85.3
+Table 5: Success rate (%) of our WorldGUI-Agent (w/ GPT-4o) with the ablation of different critical modules.
+Method Office Win. Usage Web Coding Media Overall
+Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
+Full Model 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
+– w/o Planner-Critic 31.1 17.0 20.8 9.0 38.1 25.3 36.4 11.1 20.0 10.2 18.5
+– w/o Step-Check 31.1 19.3 20.8 9.0 33.3 28.3 45.5 13.3 20.0 8.2 19.8
+– w/o Actor-Critic 15.6 7.8 4.2 3.4 28.6 17.2 0.0 8.9 10.0 6.1 9.7
+Table 6: Success rate (%) of our WorldGUI-Agent with the ablation of Instructional Video.
+Method PPT Word Excel Acrobat VSCode Overall
+Full Model 45.5 36.4 50.0 36.4 45.5 42.9
+w/o Inst. Video 45.5 27.3 25.3 18.2 27.3 28.6
+Table 7: Running time comparison.
+Method Executed Steps Time (seconds)
+Agent-S 10 131.98
+Agent-S2 9 108.65
+WorldGUI-Agent 24 129.55
+Table 8: Task counts and success rates (SR, %) across different augmentation types for WorldGUI-Agent (GPT-5.1).
+Augmentation Type #Tasks SR (%) #Success
+Add-step 255 34.1 87
+Trim-step 119 52.1 62
+Adjust-step 130 43.1 56
+Total 500 41.0 205
+[…]
+Evaluation.
+Given that our WorldGUI includes 611 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results.
+Metric.
+Following the previous works, we use Success Rate (SR) as the metric.
+Baselines.
+We implement the baseline approach called Plan-Act with different MLLMs as the base model.
+It focuses on investigating the basic capabilities of task planning and action prediction.
+Additionally, we compare our WorldGUI-Agent with two agentic frameworks and two SOTA GUI models: AssistGUI (Gao et al., 2024), Agent-S2 (Agashe et al., 2025b), Computer Use (Claud-3.5-Sonnet) (Anthropic, 2024), and UI-TARS-1.5 (Qin et al., 2025).
+AssistGUI and Agent-S2 are two prominent agentic frameworks designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query.
+We increase the base model to GPT-4o for AssistGUI and Claude-Sonnet-4 of Agent-S2 for fair performance.
+Claude Computer Use (CCU) is the leading proprietary model specially designed for computer use.
+We use the open-source implementation OOTB (Hu et al., 2024) as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison.
+We also implement our WorldGUI-Agent with four different MLLMs to illustrate the effectiveness of our proposed universal agent framework.
+6.1 Main Results on WorldGUI
+Overall.
+Table 4 reports the success rates (SR) of different agents and human experts on our WorldGUI benchmark, broken down by task type (Meta vs.
+Aug.) across five categories: Office, Win.
+Usage, Web, Coding, and Media.
+From these results we draw the following main conclusions.
+A large gap remains between agents and humans.
+The best-performing agent (WorldGUI-Agent with GPT-5.1) achieves an overall SR of only 45.8%, which is less than half of the 85.3% attained by human experts.
+This stark contrast underscores the difficulty of our tasks and the need for further advances in desktop GUI automation.
+Agents generalize poorly to augmented tasks.
+Across all methods, performance on Augmentation tasks (which introduce interface or context variations) is substantially lower than on their corresponding Meta tasks.
+For example, Claude-3.5-Sonnet in the Win.
+Usage category attains 50.0% on Meta tasks but drops to just 19.1% on Aug.
+tasks.
+This highlights the importance to evaluate the agents on various initial conditions.
+Desktop GUI tasks pose a greater challenge than web tasks.
+Every agent scores higher on Web tasks than on desktop application tasks.
+WorldGUI-Agent with GPT-5.1, for instance, jumps from 85.7% on Web Meta to only 62.5% on Win.
+Usage Meta, and the gap widens on their Augmentation counterparts.
+Thus, desktop GUI tasks remain challenges than widely-studied web GUI tasks (Cheng et al., 2024; He et al., 2024).
+WorldGUI-Agent consistently outperforms a naive Plan-Act counterpart.
+By incorporating our three critical modules into the planning and execution loop, WorldGUI-Agent substantially improves success rates over the basic Plan-Act counterpart.
+Relative to Plan-Act, WorldGUI-Agent raises overall SR by +12.2% with Gemini-2.0, +17.5% with GPT-4o, and +12.4% with Claude-3.5-Sonnet, demonstrating the effectiveness of our critical modules proposed for WorldGUI-Agent.
+Specialized GUI models transfer poorly to WorldGUI.
+Despite achieving strong results on several existing GUI benchmarks, specialized models do not perform well in our setting.
+For example, UI-TARS-1.5 attains only 12.3% overall SR on WorldGUI, which is notably lower than general-purpose MLLMs when used within our framework (e.g., 36.0% for WorldGUI-Agent with Claude-3.5-Sonnet and 45.8% with GPT-5.1), and even below the CCU baseline built on Claude-3.5-Sonnet (23.6%).
+Similarly, Agent-S2 reaches 34.2% overall SR, still far from human performance and trailing behind WorldGUI-Agent with GPT-5.1.
+These results indicate that training pipelines and architectures tuned for single-start or less diverse benchmarks do not automatically transfer to the multi-start, dynamically perturbed settings of WorldGUI.
+In our evaluation, specialized GUI models exhibit strong grounding but substantially weaker planning, which leads to low success rates even when provided with additional signals such as instructional-video subtitles.
+Table 9: Experimental results on WindowsAgentArena. The reported results are from the (Bonatti et al., 2024) and (Agashe et al., 2025b).
+Method Office Web Win. System Coding Media Win. Utils Overall
+Phi3-V (Bonatti et al., 2024) 0.0 6.9 8.3 0.0 6.2 0.0 3.5
+GPT-4o-mini (Bonatti et al., 2024) 0.0 14.9 8.3 0.0 0.0 0.0 4.2
+GPT-4o (Bonatti et al., 2024) 0.0 13.7 29.2 0.0 10.3 0.0 8.6
+NAVI (Bonatti et al., 2024) 0.0 27.3 33.3 27.3 30.3 8.3 19.5
+Agent S (Agashe et al., 2025a) w/ GPT-4o 0.0 13.3 45.8 29.2 19.1 22.2 18.2
+Agent S2 (Agashe et al., 2025b) w/ Claude-3.7-Sonnet 7.0 16.4 54.2 62.5 28.6 33.3 29.8
+WorldGUI-Agent w/ Claude-3.5-Sonnet 7.0 53.3 45.8 33.3 28.6 33.3 31.2
+6.2 Ablation Study
+Impact of different critical modules.
+Table 5 reports an ablation study of WorldGUI-Agent across five application categories.
+The full model achieves an overall success rate (SR) of 26.0%.
+Removing any core component leads to notable degradation.
+Excluding the Planner-Critic reduces SR to 18.5% (–7.5%), highlighting its role in plan refinement.
+Removing Step-Check lowers SR to 19.8% (–6.2%), mainly affecting multi-step interaction scenarios, suggesting its effectiveness in intermediate error correction.
+The most severe impact comes from removing the Actor-Critic, which collapses SR to 9.7% (–16.3%), nearly eliminating performance on Coding and Win.
+Usage tasks, underscoring the necessity of action-level reward feedback.
+Overall, the three modules provide complementary benefits, enabling robust agent performance.
+Impact of Instructional Video.
+In Table 6, we study the impact of removing the instructional video (subtitles) by modifying the prompt to include only the user query for generating the initial plan.
+In the Excel applications, we observe a significant performance decline, as their tasks are complex and difficult, and rely more heavily on additional domain knowledge for successful planning.
+In contrast, the agents performs relatively well on Win.
+Usage tasks, such as Settings and File Management, are where it has more inherent pretrained prior.
+These findings underscore the necessity of instructional videos for complex tasks like visual effect design, mirroring how users learning to build a slide often rely on tutorial videos.
+It is note that it is a setup proposed in our benchmark; one can also use our data without instructional videos.
+6.3 Results by Augmentation Type
+Across the 500 augmented tasks, WorldGUI-Agent (GPT-5.1) solves 205 cases.
+As shown in Table 8, Add-step tasks are the most challenging (34.1% SR), since inserted steps create larger state mismatches.
+Adjust-step tasks show moderate difficulty (43.1%), reflecting the need for fine-grained manipulations.
+Trim-step tasks are relatively easier (52.1%), as removing redundant steps keeps the interface closer to the meta configuration.
+These trends highlight how different forms of task-progress variation distinctively impact agent robustness.
+6.4 Failure Analysis and Robustness Dimensions
+To better understand the limitations of current GUI agents under the dynamic initial states of WorldGUI, we analyze representative failure modes drawn from our qualitative study (Appendix I).
+We observe three major robustness dimensions that frequently lead to failures.
+(1) State-mismatch robustness.
+When pre-actions introduce interface configurations different from the default view (e.g., an expanded settings dropdown or partially filled input fields), agents often misinterpret the visible state or fail to recognize that previous steps have already been executed.
+As illustrated in Figure 24 (left), once a submenu overlays other elements, the agent struggles to recover the intended target (e.g., the “System” button), leading to incorrect plan continuation or premature termination.
+(2) Fine-grained manipulation robustness.
+Tasks requiring precise adjustments—such as dragging sliders or manipulating bars to reach a specific value—remain challenging.
+Even when grounding is correct, the agent may overshoot, undershoot, or repeatedly click without achieving the required adjustment, as shown in Figure 24 (right).
+This reflects a broader difficulty in performing continuous or multi-step manipulations that depend on spatial granularity.
+(3) Ambiguous-visual-context robustness.
+In the absence of clear text cues, agents exhibit uncertainty in selecting visually similar elements.
+Figure 25 (left) illustrates a case where the model must choose a centered icon from a symmetric layout, yet struggles due to limited visual discrimination without textual anchors.
+Similarly, when multiple text elements appear close together, GUI parsers may surface misleading anchors (e.g., the label “Replace with”), causing the agent to click an irrelevant surface-level text region rather than the actual input box (Figure 25, right).
+Overall, these failure modes echo our quantitative findings: while grounding on static interfaces is often strong, robustness to state changes, fine-grained manipulations, and text-sparse visual contexts remains limited.
+These observations highlight the need for future progress in both perception and planning components, particularly under the multi-start, dynamically perturbed conditions introduced by WorldGUI.
+6.5 Computational Costs Analysis
+We select two SoTA non-training agent frameworks Agent-S (Agashe et al., 2025a) and Agent-S2 (Agashe et al., 2025b) for comparing the computational costs.
+Take a Windows setting task as an example, we tested on the same desktop PC with an AMD Ryzen 7 5800H CPU.
+As shown in Table 7, our WorldGUI-Agent shows a competitive running time of 129.55s, as compared with Agent-S (131.98s) and Agent-S2 (108.64s).
+The main computational costs of our designed modules are largely affected by calling base model.
+Since the main problem of desktop GUI automation is still the suboptimal performance as shown in Table 4, such computational costs are currently acceptable.
+There remains a clear tradeoff between performance and time costs in GUI automation, and this challenge is shared across the community.
+6.6 Results on WindowsAgentArena
+Settings.
+The official WindowsAgentArena (WAA) evaluation runs inside a Linux-based VM using a pre-built image and a JSON configuration file.
+Since the essential logic of WAA lies in its JSON specification of the environment state and task instruction, we reconstruct the same environment on a Windows machine by following the provided configuration (e.g., opening specified tabs or launching required applications).
+We then use the instruction field from the JSON file as the prompt for our WorldGUI-Agent, which executes the corresponding GUI actions on Windows.
+After completion, we follow WAA’s evaluation protocol to determine task success.
+When an official evaluation script is available, we apply it directly; otherwise, we perform manual judgment based on the given instructions.
+We do not optimize the prompts, just apply our agent directly on its tasks.
+Main results.
+Table 9 compares WorldGUI-Agent against six agents on the WAA benchmark.
+WorldGUI-Agent achieves a 31.2% overall SR surpass two SOTA api-calling agents, Agent-S and Agent-S2, even with a weaker base mode,l Claude-3.5-Sonnet.
+These results underscore the effectiveness of our designed critical modules.
+Conclusion
+We introduced WorldGUI, a benchmark that evaluates GUI agents under diverse, dynamically perturbed initial states, an aspect largely absent from existing evaluations.
+By constructing task-progress variations via pre-actions, WorldGUI enables systematic assessment of robustness and recovery in realistic GUI scenarios.
+We also presented WorldGUI-Agent, a lightweight, universal framework that inserts three verification stages into the standard planning loop to improve stability in dynamic environments.
+Our experiments show that current agents, including strong commercial models, still struggle under non-default states, while WorldGUI-Agent offers consistent gains.
+Impact Statement
+This paper introduces a benchmark and agent framework intended to advance research on dynamic GUI automation.
+Our goal is to improve the evaluation and reliability of GUI agents in controlled environments, which may support future progress in productivity tools, accessibility assistance, and educational software.
+As with many advances in machine learning and automation, there are potential risks, including misuse of automated GUI agents for large-scale or unauthorized interactions with software systems.
+Our benchmark does not involve personal or sensitive data and is designed strictly for research use.
+We encourage practitioners to incorporate appropriate safeguards, such as permission checks, sandboxing, and oversight, when deploying GUI agents in real-world applications.
+Overall, we believe the broader societal impacts of this work align with those commonly associated with progress in machine learning, and we do not identify specific concerns beyond these general considerations.
+Appendix A Task Formulation
+GUI Automation.
+Following prior work (Xie et al., 2024; Zhou et al., ), we model GUI automation as a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$.
+Given a natural-language query $q$ and optional instructional video $v$, the agent receives an observation $o_{t}$ of the current state $s_{t}$ and outputs an action $a_{t}\in\mathcal{A}$, which updates the environment to $s_{t+1}$.
+The episode terminates when the task succeeds or fails, and $\mathcal{R}$ returns a binary completion signal.
+WorldGUI Tasks.
+As illustrated in Figure 1, each task in WorldGUI is instantiated under multiple initial states that all converge to the same goal.
+We create these variants through pre-actions—short executable sequences that shift task progress without altering the UI skin, resolution, or layout.
+This design enables controlled task-progress perturbations (Add/Trim/Adjust-step) and complements existing online GUI benchmarks; key differences are summarized in Table 1.
+Observation Space.
+The agent observes (i) application metadata $m_{t}$ (e.g., panel structure and parsed UI elements), and (ii) a screenshot $V_{t}$ of the current GUI state, following  (Gao et al., 2024).
+Metadata provides structured grounding anchors, while screenshots offer holistic visual context for planning and action generation.
+Action Space.
+Actions include raw mouse and keyboard operations (click, drag, type, shortcuts), with mouse actions parameterized by pixel coordinates in $V_{t}$.
+We use PyAutoGUI [^4] https://pyautogui.readthedocs.io for execution, and represent each action in the form action_type(arguments) (Table 10).
+Table 10: The action types and examples.
+Action Type Example
+Mouse Movement moveTo(120, 200)
+Mouse Clicks click(200, 300)
+Keyboard Type write(’classes’)
+Hotkey hotkey(’ctrl’, ’a’)
+Scrolling scroll(-100)
+Drag dragTo(120, 220, 2)
+Mouse Down and Up mouseDown(); mouseUp()
+Press Keys press(’delete’)
+Key Down and Up keyDown(’shift’)
+Appendix B WorldGUI-Agent: Thinking before Doing
 Figure 6: WorldGUI-Agent.
@@
 Before each step is passed to the Actor module, it undergoes a Step-Check.
-After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes necessary corrections.
-In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  hong2024cogagent ; cheng2024seeclick ; lin2024showui ; zhang2023appagent ; AgentS .
+After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes corrections.
+In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  (Hong et al., 2024; Cheng et al., 2024; Lin et al., 2024; Zhang et al., 2023; Agashe et al., 2025a).
 The WorldGUI-Agent includes the five fundamental but essential components as in Figure 6 and an Interaction reasoning loop detailed in Algorithm 1.
 We summarize our critical designs in the following:
+Figure 7: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic provides necessary corrections.
 $\bullet$ Post-Planning Critique: After the planning phase, a critique module verifies and, if necessary, self-corrects the generated plans to ensure their accuracy.
@@
 These critique designs ensure the reliability and adaptability of WorldGUI-Agent in complex GUI environments.
-Figure 4: State-Aware Planner and Planner-Critic. The Planner generates an initial plan. Then, the Planner-Critic provides necessary corrections.
-3.1 State-Aware Planner
+B.1 State-Aware Planner
 The State-Aware Planner processes the instructional video $v$ and the user query $q$ generates an initial plan as shown in the left of Figure 7.
-We use the speech recognition model Whisper radford2023robust to translate the video $v$ into the subtitle and then send it to the MLLM for task planning.
+We use the speech recognition model Whisper (Radford et al., 2023) to translate the video $v$ into the subtitle and then send it to the MLLM for task planning.
 The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task.
-Under each $p_{i}$, there is a list of subtasks $[S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone.
+Under each $p_{i}$, there is a list of subtasks $S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone.
 To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state.
@@
 It also avoids the occlusion issue when not seeing the screenshot.
-3.2 Planner-Critic
+B.2 Planner-Critic
 Post-Planning Critique.
@@
 (4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons.
-As related works, CoT wei2022chain , OpenAI-o1 openaio1 , and Deepseek-R1 deepseekr1 demonstrate that generating reasoning steps along with the answer would enhance the performance.
-3.3 Step-Check
+As previous studies (e.g., CoT (Wei et al., 2022), Deepseek-R1 (DeepSeek-AI et al., 2025)) demonstrate that generating reasoning steps along with the answer would enhance the performance.
+B.3 Step-Check
+Figure 8: Step-Check. This module first checks the step completion status via an MLLM and then navigates to the current task processing.
 Pre-Action Validation.
@@
 Therefore, the model must determine whether to proceed with a subtask $S_{t}$ based on the current state (screenshot: $V_{t}$, metadata: $M_{t}$).
-As illustrated in Figure 5, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into four types:
+As illustrated in Figure 8, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into:
 (1) <Modify>: Indicates that the subtask should be modified or additional subtasks should be added.
 (2) <Pass>: Indicates that the current subtask is unnecessary and can be skipped.
-Figure 5: Step-Check. This module first checks the step completion status via an MLLM and then navigates current task processing.
 (3) <Continue>: Indicates that the subtask is valid and should be executed as planned.
 (4) <Finished>: Indicates that the subtask has already been completed and requires no further action.
-In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs ##\##Cannot confirm”.
+In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs #\#Cannot confirm”.
 When this occurs, we design a Region Search module implemented by an LLM.
@@
 The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
-3.4 Actor
+B.4 Actor
 The goal of the Actor is to translate natural language subtask $S_{t}$ into executable code $C_{t}$.
@@
 The generated actions will be executed in the environment, and then the new screenshot $V_{t+1}$ and metadata $m_{t+1}$ will be captured for the next processing.
-3.5 Actor-Critic
+B.5 Actor-Critic
 Figure 9: Actor-Critic. This module includes two parts: task verification and task correction. The design follows the verify-then-correct mechanism.
 […]
-4 Experiments
-Table 4: Success rate (%) of different agents on WorldGUI.
-Human denotes the average performance of four expert participants who have watched the instructional video only once, similar to the model.
-Meta represents the meta task, while Aug.
-represents the augmented task.
-Method MLLM Office Win. Usage Web Coding Media Overall
+Table 11: Performance comparison between WorldGUI-Agent (with Claude-3.5-Sonnet and Claude-Sonnet-4) and Agent-S2 (Claude-Sonnet-4). Results are reported across five representative applications.
+Method PPT VSCode Acrobat VLC File Explorer
 Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
-Plan-Act Gemini-2.0 8.9 3.2 8.3 3.4 28.6 16.2 18.2 2.2 10.0 2.0 6.9
-Plan-Act GPT-4o 13.3 10.1 8.3 2.3 23.8 11.1 9.1 2.2 10.0 2.0 8.5
-AssistGUI GPT-4o 26.7 16.1 29.2 7.9 33.3 20.2 27.3 11.1 10.0 8.2 16.5
-Computer Use Claude-3.5 28.9 19.3 29.2 14.6 71.4 32.3 54.6 22.2 30.0 6.1 23.6
-WorldGUI-Agent Gemini-2.0 31.1 17.0 20.8 9.0 38.1 29.3 36.4 11.1 20.0 10.2 19.1
-WorldGUI-Agent GPT-4o 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
-WorldGUI-Agent Claude-3.5 57.8 32.6 50.0 19.1 76.2 46.5 54.6 26.7 50.0 18.4 36.0
-Human 88.9 83.5 100.0 89.9 95.2 80.8 81.8 77.8 90.0 85.7 85.3
-Implementation Details.
-We implement the MLLM in our WorldGUI-Agent by using GPT-4o gpt4o (gpt-4o-2024-08-06) by default.
-For the computer mouse and keyboard control, we use the Python library PyAutoGUI.
-Following the AssistGUI assistgui , we use the GUI parser to obtain the position information of elements, e.g., buttons, icons, and text.
-We use some vision foundation models, such as Google OCR, to extract the text.
-By default, we use the center coordinates to represent the location of each element.
-All the testing is under the same screenshot resolution (1920 $\times$ 1080).
-In all experiments, we set the max trials of the Actor-Critic to 3 for light interaction costs.
-For the total trials of each task, we set it to $14\times N+$ 1, where $N$ is set empirically.
-Evaluation.
-Given that our WorldGUI includes 611 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results.
-Metric.
-Following the previous works of OSworld and AssistGUI, we use Success Rate (SR) as the metric.
-Baselines.
-We implement the baseline approach called Plan-Act with different MLLMs as the base model.
-It focuses on investigating the basic capabilities of task planning and action prediction.
-Additionally, we compare our WorldGUI-Agent with two strong approaches: AssistGUI assistgui and Computer Use (Claude 3.5) claude3.5 .
-AssistGUI is a prominent agent framework designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query.
-We implement it by increasing the MLLM to GPT-4o for better performance.
-Computer Use (Claude 3.5) is the leading proprietary model specially designed for autonomous computer use.
-We use the open-source implementation OOTB hu2024dawnguiagentpreliminary as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison.
-We also implement our WorldGUI-Agent with three different MLLMs to illustrate the effectiveness of our proposed universal agent framework.
-4.1 Main Results on WorldGUI
-Table 4 reports the success rates (SR) of different agents and human experts on our WorldGUI benchmark, broken down by task type (Meta vs.
-Aug.) across five categories: Office, Win.
-Usage, Web, Coding, and Media.
-From these results we draw the following main conclusions.
-A large gap remains between agents and humans.
-The best-performing agent (WorldGUI-Agent with Claude-3.5) achieves an overall SR of only 36.0%, which is less than half of the 85.3% attained by human experts.
-This stark contrast underscores the difficulty of our tasks and the need for further advances in desktop GUI automation.
-Agents generalize poorly to augmented tasks.
-Across all methods, performance on Augmentation tasks (which introduce interface or context variations) is substantially lower than on their corresponding Meta tasks.
-For example, Claude-3.5 in the Win.
-Usage category attains 50.0% on Meta tasks but drops to just 19.1% on Aug.
-tasks.
-This highlights the importance of dynamic testing to capture realistic human–computer interaction.
-Desktop applications pose a greater challenge than web tasks.
-Every agent scores higher on Web tasks than on desktop application tasks.
-WorldGUI-Agent with Claude-3.5, for instance, jumps from 76.2% on Web Meta to only 57.8% on Office Meta, and the gap widens on their Augmentation counterparts.
-Thus, desktop GUI automation remain a frontier for computer use research.
-WorldGUI-Agent consistently outperforms a naive Plan-Act baseline.
-By incorporating our three critical modules into the planning and execution loop, WorldGUI-Agent substantially improves success rates over the basic Plan-Act approach.
-Relative to Plan-Act, WorldGUI-Agent raises overall SR by +12.2% with Gemini-2.0, +17.5% with GPT-4o, and +12.4% with Claude-3.5, demonstrating the effectiveness of our design across multiple MLLMs.
-Table 5: Success rate (%) of our WorldGUI-Agent (Full Model) with the ablation of different critical modules.
-Method Office Win. Usage Web Coding Media Overall
-Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
-Full Model 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
-– w/o Planner-Critic 31.1 17.0 20.8 9.0 38.1 25.3 36.4 11.1 20.0 10.2 18.5
-– w/o Step-Check 31.1 19.3 20.8 9.0 33.3 28.3 45.5 13.3 20.0 8.2 19.8
-– w/o Actor-Critic 15.6 7.8 4.2 3.4 28.6 17.2 0.0 8.9 10.0 6.1 9.7
-Table 6: Success rate (%) of our WorldGUI-Agent with the ablation of Instructional Video (Inst. Video).
-Method PPT Word Excel Acrobat VSCode Overall
-Full Model 45.5 36.4 50.0 36.4 45.5 42.9
-w/o Inst. Video 45.5 27.3 25.3 18.2 27.3 28.6
-4.2 Ablation Study
-Impact of different critical modules.
-Table 5 presents the results of an ablation study on the three core components of WorldGUI-Agent across five application categories (Office, Windows Usage, Web, Coding, Media).
-The full model achieves an overall success rate (SR) of 26.0%.
-The effects of removing each component are as follows: Planner-Critic: Eliminating this module reduces overall SR to 18.5% (–7.5%), with substantial drops in Office (42.2% → 31.1%) and Web (47.6% → 38.1%) tasks, indicating its importance for refining initial plans.
-Step-Check: Without step-wise verification, SR decreases to 19.8% (–6.2%).
-The relatively smaller decline on Coding and Win.
-Usage tasks suggest that Step-Check excels at intercepting and correcting multi-step interaction errors.
-Actor-Critic: Removing the action-level critic causes SR to collapse to 9.7% (–16.3%).
-Performance on Coding Meta drops to 0.0% and Windows Usage Meta to 4.2%, highlighting the critical role of reward-driven action correction for action-level GUI operations.
-These results confirm that Planner-Critic, Step-Check, and Actor-Critic each contribute complementary benefits—plan refinement, intermediate validation, and action optimization—that are essential for the robustness and overall effectiveness of WorldGUI-Agent.
-Impact of Instructional Video.
-In Table 6, we study the impact of removing the instructional video by modifying the prompt to include only the user query for generating the initial plan.
-In the Excel applications, we observe a significant performance decline, as their tasks are complex and difficult, and rely more heavily on additional domain knowledge for successful planning.
-In contrast, the MLLM performs relatively well on Win.
-Usage tasks, such as Settings and File Management, are where it has more inherent familiarity.
-These findings underscore the necessity of instructional videos for complex tasks like visual effect design, mirroring how users learning to build a slide often rely on tutorial videos.
-4.3 Results on WindowsAgentArena Benchmark
-Table 7 compares WorldGUI-Agent against four leading agents on the WindowsAgentArena benchmark.
-WorldGUI-Agent achieves a 30.5% overall SR, far surpassing GPT-4V (19.5%), GPT-4o (8.6%), GPT-4o-mini (4.2%), and Phi3-V (3.5%).
-Its gains are most pronounced in desktop categories: Office tasks (4.7% vs.
-0%), Windows System (45.8% vs.
-33.3%), and Windows Utilities (33.3% vs.
-8.3%).
-On web browser, it reaches 53.3%, nearly double GPT-4V’s 27.3%, and on coding tasks, it records 33.3% versus 27.3%.
-In media tasks, WorldGUI-Agent posts 28.6%, closely matching GPT-4V’s 30.3%.
-These results underscore the necessity of integrated planning critique, step-check verification, and action-level feedback.
-These results demonstrate that our framework robustly handles both desktop GUI tasks and dynamic web environments, highlighting its versatility for real-world GUI automation.
-Table 7: Experimental results on WindowsAgentArena windowsagentarena . The reported results are from the windowsagentarena .
-Method Office Web Win. System Coding Media Win. Utils Overall
-Phi3-V 0.0 6.9 8.3 0.0 6.2 0.0 3.5
-GPT-4o-mini 0.0 14.9 8.3 0.0 0.0 0.0 4.2
-GPT-4o 0.0 13.7 29.2 0.0 10.3 0.0 8.6
-GPT-4V 0.0 27.3 33.3 27.3 30.3 8.3 19.5
-WorldGUI-Agent (Ours) 7.0 53.3 45.8 33.3 28.6 33.3 31.2
-Conclusion
-In this paper, we take the first step toward comprehensive GUI agent evaluation by introducing WorldGUI.
-In addition to the standard static testing processes, we incorporate dynamic testing procedures to ensure that WorldGUI effectively captures the complexity and dynamism of real-world GUI environments.
-Furthermore, we propose a universal agent framework, WorldGUI-Agent, built upon the critical thinking principle.
-This framework enables the agent to dynamically identify uncommon states and adjust its plans or actions accordingly.
-Finally, we evaluate WorldGUI-Agent powered by Claude-3.5-Sonnet on WorldGUI and WindowsAgentArena benchmarks, demonstrating the effectiveness across a variety of GUI tasks.
-Limitation and Society Impacts
-In the current implementation of our agent, WorldGUI-Agent, more external tools have not been integrated into the GUI planning and action prediction processes to prioritize computational efficiency.
-Incorporating tools such as web search or file search into the agent’s design could be a valuable future direction to improve performance.
-Additionally, due to the usage of the GUI Parser, which would increase the time costs because of the response speed of experimental desktop computers, it is still a tradeoff between performance and running time in the current GUI domain.
-We consider that if the base MLLM model is specifically trained with stronger planning and grounding ability, the running time would be sped up.
-It is noted that our agent framework is capable of working with any MLLM.
-WorldGUI takes the first step of pushing the GUI automation into the dynamic testing process, as we found that real-world human-computer interactions are dynamic and unpredictable; existing GUI benchmarks fail to capture such dynamics to closely reflect the interactions.
-Our WorldGUI-Agent is a straightforward and universal agent framework by considering incorporates three critical modules to adaptively align the plan and actions with exact environment situations, which would be a good baseline for future agent development.
-For instance, incorporating more tools such as web search or file search into the planning module or action prediction module to realize more challenging tasks.
-Appendix A Related Work
-A.1 GUI Benchmarks
-GUI benchmarks are essential for evaluating the performance and robustness of GUI agents.
-For web applications, WebShop [29], WebArena [34], and WebVoyager [10] focus on creating the GUI tasks in a web browsing scenario.
-In OS environments, OSWorld [28] is a comprehensive benchmark, including various operating systems with real applications.
-Mobile benchmarks such as MobileAgent [23] and AppAgent [31] propose two GUI benchmarks of mobile applications.
-Windows-related benchmarks like AssistGUI [7] and WindowAgentArena [4] propose a list of real tasks in the Windows platform.
-However, these online testing GUI benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments.
-As a result, they are insufficient for comprehensively evaluating GUI agents.
-A.2 GUI Agents
-CogAgent [12] is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick [5] and SeeAct [32] focus on the GUI grounding for enhancing the task performance.
-MobileAgent [23] and AppAgent [31] are proposed to design the agent on the mobile device.
-Ferret-UI [30] is another representative work focusing on enhancing the grounding ability in the IOS platform.
-These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction, but still face limitations in handling dynamic and complicated full GUI tasks.
-Therefore, to enhance GUI automation in dynamic environments, we propose WorldGUI-Agent, which improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes.
-The components comparison of our WorldGUI-Agent and other closely related agents is shown in Table 8.
-Table 8: Comparison with other closely related agents.
-Most existing agents solely focus on post-action evaluation but omit the post-planning critique and pre-action validation in handling dynamic GUI environments.
-Method Post-Planning Critique Pre-Action Validation Post-Action Evaluation
-Mobile-Agent [23]
-Mobile-Agent-V2 [23]
-AssistGUI [7]
-Agent-S [1]
-Mobile-Agent-E [24]
-WorldGUI-Agent (ours)
-A.3 Critical Thinking in Agents
-Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 [19] and Deepseek-R1 [6], have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks.
-The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency.
-One notable LLM-based agent framework, Reflexion [22], demonstrates the effectiveness of self-reflection in solving complex tasks.
-Furthermore, CRITIC [8] integrates external tools into the critique process, leveraging them to improve performance.
-Noticing that the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario.
-Which is not only aims to enhance the reasoning performance but is also indispensable to designing the key module Actor-Critic [15] to ensure task completion.
-A closely related work, AssistGUI [7], integrates a critical module only after the Actor module to evaluate action completion.
-Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor.
-These two modules lead to a universal and fundamental GUI agent framework WorldGUI-Agent which will provide insights for future GUI agent design.
-Appendix B Data
-Annotators.
+Agent-S2 w/ Claude-Sonnet-4 45.5 18.9 45.5 37.8 18.2 14.5 20.0 16.3 60.0 64.7
+WorldGUI-Agent w/ Claude-3.5-Sonnet 54.5 39.6 54.5 26.7 63.6 20.0 50.0 18.4 50.0 17.6
+WorldGUI-Agent w/ Claude-Sonnet-4 63.6 52.8 54.5 28.9 54.5 30.9 40.0 28.6 70.0 26.5
+Table 12: Performance comparison between WorldGUI-Agent (with UI-TARS-1.5) and UI-TARS-1.5.
+Method PPT Acrobat
+Meta Aug. Meta Aug.
+WorldGUI-Agent w/ UI-TARS-1.5 36.6 18.9 36.4 9.1
+UI-TARS-1.5 27.3 17.0 9.1 1.8
+Appendix C Additional Experiments
+As shown in Table 11, Agent-S2 (Agashe et al., 2025b) shows competitive results as compared with our WorldGUI-Agent.
+We also test on two representative office software to compare the effectiveness of our proposed agentic framework by replacing the base model with UI-TARS-1.5 (Qin et al., 2025) in Table 12.
+It is noted that to improve the performance of UI-TARS-1.5, we use the GPT-4o to task planning, as we found that UI-TARS struggles with understanding complex desktop software layout and cannot capture the dynamic initial condition changes.
+We use GPT-4o for better implementation.
+Appendix D Data
+D.1 Annotators
 In this work, we have four annotators: A, B, C, and D.
@@
 Agent: We design an agent built upon the MLLMs to achieve the goal.
-Creating Augmented Tasks.
-In our study, to stimulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks.
+D.2 Creating Augmented Tasks
+In our study, to simulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks.
 Specifically, we propose pre-actions before executing the task.
@@
 If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
-Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings.
+D.3 Introducing Diverse Initial Context States
+Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings.
 This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps.
@@
 We additionally provide more examples about augmenting the meta task in Figure 11.
-Figure 8: An example of augmenting one GUI task with manually aug the initial state and then using the execution scripts and corresponding agents to obtain the pre-action for each augmented case.
 Figure 11: We present the examples of conducting the augmentations from the meta task.
-Appendix C Data Statistics
-Augmentation tasks type analysis.
-As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state.
-(2) The human-computer interactions may start from the intermediate state of a specific task.
-We propose to create the augmented tasks for each meta task to stimulate authentic GUI interactions.
-Our augmentations lie in two main groups: (1) stimulating the intermediate states and (2) introducing diverse initial states.
-We divide the two groups into three exact types: Add-step, Trim-step, and Adjust-step.
-For Add-step, it represents various unrelated state augmentations to stimulate the scenario that we may start the agent-computer interactions in another unrelated task or interfaces, the agent should replan the task to add necessary steps.
-For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state.
-For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu.
-Most of the time, it would not require new steps to return to the target task progress.
-This augmentation may mislead the agent in state understanding, making them jump or miss the key steps.
-As shown in Figure 11, the manually created augmentations mainly belong to the add-step.
-Adjust-step could be the second-largest application, except for the File Explorer.
-Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
-Task difficulty analysis.
-Figure 12 shows the distribution of the task difficulty across desktop applications.
-We annotated the task difficulty level based the subjective software usage experience.
-The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging.
-The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels.
-By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents.
-According to VSCode and File Explorer, and YouTube applications, the tasks are easier than in other applications.
-Overall, the task difficulty of created data is diverse across different applications, and there is still a need for stronger agents focusing on handling desktop-oriented GUI tasks.
-The task details about the user query and the pre-actions are included in the metadata JSON file in the supplementary materials.
-The project file, instruction video, and augmentation files can be found in the provided data link.
-Figure 10: Distribution of Software taxonomy and the distribution of task length.
-Figure 11: The distribution of different augmentation types.
-Figure 12: The distribution of different task difficulty.
-Table 10: Task statistics of WorldGUI.
-Category Software Meta Count Aug Count Meta-Task Length (Avg.)
-Office Adobe Acrobat 11 55 7.1
-Office Excel 12 58 4.1
-Office PowerPoint 11 53 5.1
-Office Word 11 52 5.1
-Coding VSCode 11 45 4.4
-Windows Usage File Explorer 10 34 7.1
-Windows Usage Settings 14 55 5.1
-Web Usage Web Browser 10 49 7.8
-Web Usage YouTube 11 50 4.8
-Media VLC Player 10 49 10.8
-Total (Average) 111 500 6.0
-Appendix D Detailed Experimental Results
-Table 11 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena [4] benchmark.
+Appendix E Detailed Experimental Results
+Table 14 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024) benchmark.
 The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks.
 As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
-Table 11: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena [4].
+Table 14: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024).
 Domain Application #Tasks #Successes SR (%)
@@
 Overall 154 48 31.17
-Appendix E Computational Costs Discussion
+Appendix F Computational Costs Discussion
 The average number of execution steps and tokens consumed are shown below Table 15.
@@
 Web $\sim$17 1637 84
-Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a Laptop with AMD Ryzen 7 4800H CPU in Table 13.
+Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a desktop PC with AMD Ryzen 7 5800H CPU.
 Task length: 6 (generated by Planner+Planner-Critic).
-Total steps: 18, total time: 335.84s, total agent execution time: 412.37s.
-Since desktop GUI is in its early stages, computational costs are currently unavoidable.
-Even OpenAI’s Deep Research takes over 10 minutes in daily use.
-According to the post from OpenAI’s Operator, achieving 38.1% on OS-World requires over 100 steps, which is costly as well.
-In summary, it is still a tradeoff between performance and time costs in GUI automation.
-Table 13: Time costs of each module.
+To facilitate a fair comparison, we additionally selected two of the latest SOTA agents, Agent-S (2024-10-08) and Agent-S2 (2025-04-01), and measured their run times on the same successful task under identical hardware and the same base MLLM (Claude-Sonnet-4).
+The results are shown in Tables 16, 17, 18.
+To summarize, our WorldGUI-Agent shows a competitive running time of 129.55s, as compared with Agent-S (131.98s) and Agent-S2 (108.64s).
+The main computational costs of our designed modules are largely affected by the underlying large multimodal model, leaving room for acceleration optimization.
+Since desktop GUI automation is still in its early stages, such computational costs are currently unavoidable.
+For reference, even OpenAI’s Deep Research reportedly takes over 10 minutes in daily usage.
+According to OpenAI Operator’s report, achieving 38.1% on OS-World requires over 100 steps, which is similarly costly.
+In summary, there remains a clear tradeoff between performance and time costs in GUI automation, and this challenge is shared across the community.
+Table 16: Running time with WorldGUI-Agent (ours).
 Subtask Index Executed Modules Time (seconds)
-0 Planner+Planner-Critic 12.53
-1 Step-Check+Actor 21.23
-1 Actor-Critic 13.07
-2 Step-Check+Actor 17.81
-2 Actor-Critic 11.86
-3 Step-Check+Actor 18.43
-3 Actor-Critic 11.72
-4 Step-Check+Actor 22.49
-4 Actor-Critic 25.31
-4 Actor-Critic 10.82
-5 Step-Check+Actor 17.01
-5 Actor-Critic 22.36
-5 Actor-Critic 26.85
-5 Actor-Critic 18.99
-6 Step-Check+Actor 21.79
-6 Actor-Critic 23.37
-6 Actor-Critic 21.12
-6 Actor-Critic 19.08
-Appendix F Examples of Augmentations
-In this section, we present several augmentation examples in Figures 13, 14, 15, 16, 17, 18, 19.
+0 Planner 4.48
+0 Planner-Critic 11.47
+1 Parser 2.03
+1 Step-Check 4.95
+1 Actor 7.47
+1 Parser 2.07
+1 Actor-Critic 7.38
+2 Parser 2.03
+2 Step-Check 6.71
+2 Actor 6.05
+2 Parser 2.07
+2 Actor-Critic 7.97
+3 Parser 2.12
+3 Step-Check 6.35
+3 Actor 6.71
+3 Parser 2.22
+3 Actor-Critic 10.06
+4 Parser 2.01
+4 Step-Check 6.19
+4 Actor 8.54
+4 Parser 2.40
+4 Actor-Critic 9.55
+5 Parser 2.22
+5 Step-Check 6.50
+Total - 129.55
+Average (per action) - 22.72
+Table 17: Running time with Agent-S.
+Step Executed Modules Time (seconds)
+0 Manager 2.02
+1 Manager 10.11
+2 Worker 16.05
+3 Worker 12.38
+4 Worker 17.30
+5 Worker 12.26
+6 Worker 15.17
+7 Worker 15.46
+8 Worker 11.83
+9 Worker 19.40
+Total - 131.98
+Average (per worker) - 14.98
+Table 18: Running time with Agent-S2.
+Step Executed Modules Time (seconds)
+0 Manager 11.87
+1 Worker 7.27
+2 Worker 13.21
+3 Worker 13.59
+4 Worker 14.04
+5 Manager 10.61
+6 Worker 9.87
+7 Manager 7.53
+8 Worker 20.65
+Total - 108.64
+Average (per worker) - 13.22
+Appendix G Examples of Augmentations
+In this section, we present several augmentation examples in Figures 12, 13, 14, 15, 16, 17, 18.
 It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step.
@@
 Figure 18: Augmented example of a PowerPoint Task.
-Appendix G WorldGUI-Agent Reasoning Loop Algorithm
+Appendix H WorldGUI-Agent Reasoning Loop Algorithm
 In this section, we provide the details of our reasoning loop algorithm in Algorithm 1.
@@
  Generate task plan $p$ with Planner and Planner-Critic
-  Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of $p$.
-  Initial $s_{0}=\textless Continue\textgreater$ >
+ Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of pp.
+ Initial $s_{0}=\textless Continue\textgreater$
 while $S_{t}$ is not end and $t<$ max trials do
@@
   Obtain state $s_{t}$ by running Step-Check.
-     if $s_{t}=\textless Next\textgreater$ > then
+  if $s_{t}=\textless Next\textgreater$ then
    Go to the next task $S_{t+1}=next(S_{t})$
@@
   Obtain action code $C_{t}$ by running Actor; Execute the action code $C_{t}$ in the Env.; Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
-     Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ > (For each subtask, the first step is finished, then execute the actor-critic process)
+  Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ (For each subtask, the first step is finished, then execute the actor-critic process)
   Observe metadata $m_{t}$ and Screenshot $V_{t}$ from Env.
   Running Actor-Critic and obtain the state $s_{t}$
-     if $s_{t}=\textless Next\textgreater$ > then
+  if $s_{t}=\textless Next\textgreater$ then
    Go to the next task $S_{t+1}=next(S_{t})$.
   end if
-     while $s_{t}=\textless Critic\textgreater$ > and $z<$ max critique trials do
+  while $s_{t}=\textless Critic\textgreater$ and $z<$ max critique trials do
    Running Actor-Critic and obtain the state $s_{t}$ and corrected action code $C_{t}$
-        if $s_{t}=\textless Next\textgreater$ > then
+   if $s_{t}=\textless Next\textgreater$ then
     Go to the next task $S_{t+1}=next(S_{t})$.
@@
 end while
-Appendix H Qualitative Results
+Appendix I Qualitative Results
 (1) In Figure 19, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code.
 Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
-Figure 20: We show one successful prediction of our WorldGUI-Agent.
-Figure 21: We show two examples of using GUI Parser to obtain the element position information.
 (2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 21, Figure 22, and Figure 23.
 These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
-Figure 22: An example of using Planner-Critic to correct the plan.
-Figure 23: Two examples of using Step-Check to check the subtask status.
-Figure 24: An example of using Actor-Critic to correct the actions.
 (3) We also highlight some common errors encountered.
-1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application.
-As shown in the left of Figure 25, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden.
+1) The model has difficulty obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application.
+As shown on the left of Figure 24, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden.
 Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area.
@@
 The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot.
-4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task success.
+4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task’s success.
+Figure 19: We show one successful prediction of our WorldGUI-Agent.
+Figure 20: We show two examples of using GUI Parser to obtain the element position information.
+Figure 21: An example of using Planner-Critic to correct the plan.
+Figure 22: Two examples of using Step-Check to check the subtask status.
+Figure 23: An example of using Actor-Critic to correct the actions.
 Figure 24: We display some common errors.